### v2

# Asymptotic stability of composite waves of two viscous shocks for relaxed compressible Navier-Stokes equations

Abstract.
This paper investigates the time asymptotic stability of composite waves formed by two shock waves within the context of one-dimensional relaxed compressible Navier-Stokes equations. We establish the nonlinear stability of the composite waves consisting of two viscous shocks under the condition of having two small, independent wave strengths and the presence of small initial perturbations. Furthermore, the solutions of the relaxed system are observed to globally converge over time to those of the classical system as the relaxation parameter approaches zero. The methods are based on relative entropy, the $a$-contraction with shifts theory and fundamental energy estimates.
Keywords: Relaxed compressible Navier-Stokes equations; asymptotic stability; relaxation limit; composite waves of two shock wave; relative entropy; energy estimates
AMS classification code: 76N15, 35B35, 35Q35
Renyong Guan, Department of Mathematics, China University of Mining and Technology, Beijing, 100083, P.R. China, renyguan@163.com
Yuxi Hu, Department of Mathematics, China University of Mining and Technology, Beijing, 100083, P.R. China, yxhu86@163.com
1. Introduction
In this paper, we study the one-dimensional isentropic compressible Navier-Stokes equations, complemented by Maxwell’s constitutive relations. The equations are formulated as follows:
$\begin{cases}\rho_{t}+(\rho u)_{x}=0,\\ (\rho u)_{t}+(\rho u^{2})_{x}+p(\rho)_{x}=\Pi_{x},\\ \tilde{\tau}(\rho)(\Pi_{t}+u\Pi_{x})+\Pi=\mu u_{x},\end{cases}$ (1.1)
where $(t,x)\in(0,+\infty)\times\mathbb{R}$. Here, $\rho$, $u$, $\Pi$ represent fluid density, velocity and stress, respectively. $\mu>0$ is the viscosity constant. The pressure $p$ is assumed to satisfy the usual $\gamma$-law, $p(\rho)=A\rho^{\gamma}$ where $\gamma>1$ denotes the adiabatic index and $A$ is any positive constant. Without loss of generality, we assume $A=1$ in the sequel.
The constitutive relation (1.1)3italic-(1.1subscriptitalic-)$_{3}$, first introduced by Maxwell in [35], serves to delineate the connection between the stress tensor and velocity gradient for non-simple fluid. The relaxation parameter $\tilde{\tau}=\tilde{\tau}(\rho)$ describe the time lag in response of the stress tensor to velocity gradient. In fact, even for simple fluid, water for example, the time lag does exists but it is very small ranging from 1 ps to 1 ns, see [34, 40]. However, Pelton et al. [38] showed that such a time lag cannot be neglected, even for simple fluids, in the experiments of high-frequency (20GHZ) vibration of nano-scale mechanical devices immersed in water-glycerol mixtures. It was shown that, see also [2], equation (1.1)3italic-(1.1subscriptitalic-)$_{3}$ provides a general formalism with which to characterize the fluid-structure interaction of nano-scale mechanical devies vibrating in simple fluids.
Assuming that $\tilde{\tau}(\rho)=\tau\rho$, where $\tau$ is a positive constant, equation (1.1)3italic-(1.1subscriptitalic-)$_{3}$ reduces to
$\tau\rho(\Pi_{t}+u\Pi_{x})+\Pi=\mu u_{x}.$ (1.2)
[…]
$\begin{cases}v_{t}-u_{x}=0,\\ u_{t}+p_{x}=\Pi_{x},\\ \tau\Pi_{t}+v\Pi=\mu u_{x},\end{cases}$ (1.3)
where $v=\frac{1}{\rho}$ denotes the specific volume per unit mass.
We are interested in the Cauchy problem to system (1.3) for the functions
$(v,u,\Pi):[0,+\infty)\times\mathbb{R}\rightarrow(0,\infty)\times\mathbb{R}\times\mathbb{R}$
with initial conditions
$(v,u,\Pi)(0,x)=(v_{0},u_{0},\Pi_{0})(x)\rightarrow(v_{\pm},u_{\pm},0)\quad(x\rightarrow\pm\infty).$ (1.4)
The large-time behavior of solutions to system (1.3)-(1.4) is closely related to the Riemann problem of the associated $p$-system
$\begin{cases}v_{t}-u_{x}=0,\\ u_{t}+p(v)_{x}=0,\end{cases}$ (1.5)
with the Riemann initial data:
$(v,u)(t=0,x)=\begin{cases}(v_{-},u_{-}),\quad x<0,\\ (v_{+},u_{+}),\quad x>0.\end{cases}$ (1.6)
We recall that there exists a unique intermediate state $(v_{m},u_{m})$ connected to $(v_{-},u_{-})$ and $(v_{+},u_{+})$ by 1-shock wave and 2-shock wave, respectively. And it satisfies Rankine-Hugoniot condition
$\begin{cases}\sigma_{1}(v_{m}-v_{-})=-(u_{m}-u_{-}),\\ \sigma_{1}(u_{m}-u_{-})=p(v_{m})-p(v_{-}),\end{cases}\begin{cases}\sigma_{2}(v_{+}-v_{m})=-(u_{+}-u_{m}),\\ \sigma_{2}(u_{+}-u_{m})=p(v_{+})-p(v_{m}),\end{cases}$ (1.7)
and Lax entropy condition
$-\sqrt{-p^{\prime}(v_{-})}<\sigma_{1}<-\sqrt{-p^{\prime}(v_{m})},$ (1.8)
$\sqrt{-p^{\prime}(v_{m})}<\sigma_{2}<\sqrt{-p^{\prime}(v_{+})}.$
Then, the Riemann problem of the associated $p$-system has a composite wave solution $(\bar{v},\bar{u}):=(v_{1}^{s},u_{1}^{s})+(v_{2}^{s},u_{2}^{s})-(v_{m},u_{m})$, where 1-shock wave solution $(v_{1}^{s},u_{1}^{s})$ and 2-shock wave solution $(v_{2}^{s},u_{2}^{s})$ defined as follows (see [41]):
$(v_{1}^{s},u_{1}^{s})(t,x)=\begin{cases}(v_{-},u_{-}),\quad x<\sigma_{1}t,\\ (v_{m},u_{m}),\quad x>\sigma_{1}t,\end{cases}(v_{2}^{s},u_{2}^{s})(t,x)=\begin{cases}(v_{m},u_{m}),\quad x<\sigma_{2}t,\\ (v_{+},u_{+}),\quad x>\sigma_{2}t.\end{cases}$
If $\tau=0$, the system (1.3) reduces to classical compressible isentropic Navier-Stokes equations:
$\begin{cases}v_{t}-u_{x}=0,\\ u_{t}+p(v)_{x}=\left(\mu\frac{u_{x}}{v}\right)_{x}.\end{cases}$ (1.9)
The asymptotic behavior of solutions for system (1.9) and its non-isentropic counterpart has been extensively explored with a variety of initial conditions, as documented in [9, 10, 15, 16, 25, 28, 29, 30]. Notably, for shock profile initial data, Matsumura and Nishihara [28], Goodman [7] pioneered the establishment of the stability of traveling waves with sufficiently small initial disturbances and zero mass condition, by utilizing the anti-derivative method. Furthermore, Huang and Matsumura [16] demonstrated the asymptotic stability of a composite wave consisting of two viscous shocks within the Navier-Stokes-Fourier system, given that the shocks are of small magnitude and of the same order.
In a recent advancement, Kang, Vasseur and Wang [25] overcame the challenge of reconciling the standard anti-derivative method typically employed for viscous shock stability with the energy method used for rarefaction stability. They affirmed the stability of composite waves comprising both viscous shock and rarefaction by employing the method of relative entropy and the $a$-contraction with shifts theory. These methods was initially introduced by Bresch and Desjardins in [1] and further developed by Kang and Vasseur in [18, 19, 20, 23], with additional insights provided in [21, 24, 27, 44]. Employing similar methodologies, Han, Kang and Kim [10] have recently established the uniform convergence toward a composite of two viscous shocks for system (1.9) with independently small amplitudes. The objective of the present paper is to extend the results in [10] to the context of the relaxed compressible Navier-Stokes equations (1.9).
For relaxed compressible Navier-Stokes equations, the time-asymptotic stability of both single viscous shock waves and composites of two rarefaction waves has been a subject of study. When $\tilde{\tau}(\rho)$ is a constant, Hu-Wang [13] and Hu-Wang [14] respectively established the linear stability of the viscous shock wave and nonlinear stability of rarefaction waves. With $\tilde{\tau}=\tau\rho$, by checking Majda’s condition on the Lopatinski determinant and Zumbrun’s Evans function condition [31, 32, 37], Freistühler [5] get the nonlinear stability of the viscous shock waves for system (1.1) with shock profile initial data. More recently, the authors [8] have shown the nonlinear stability of composite waves of viscous shock and rarefaction, where the method of relative entropy and the $a$-contraction with shifts theory were fully used.
In this paper, we studied the time asymptotic stability for composite waves of two viscous shocks to system (1.3)-(1.4). Note that the dissipation structure of relaxed system (1.3) is much weaker than that of classical system (1.9), thus the BD entropy used in [25, 10] to prove the $a$-contraction property is not available for our system and energy estimates have new challenges. Moreover, compared to the results of [10] where cutoff functions were used to localize the diffusion term, we do not need this kind of operation. We shall follow the methods in [8] to overcome such difficulties. Here are our strategy. Instead of using BD entropy, we use of the special hyperbolic structure of the relaxed system and the relative entropy quantities with weight function and shifts to get the $L^{\infty}_{t}L^{2}_{x}$ estimates of $(v-\widetilde{v},u-\widetilde{u},\Pi-\widetilde{\Pi})$ and weighted $L^{2}_{t}L^{2}_{x}$ estimates of $(v-\widetilde{v},\Pi-\widetilde{\Pi})$ (see Lemma 4.2). We note that, unlike that in [25, 10] where the dissipation estimates of the derivative of $(v-\widetilde{v})$ were obtained for system (1.3), we do not have such estimate. Instead, we shall recover the dissipation estimate of $(v-\widetilde{v})$ from high-order estimates and thus cutoff functions are not needed.
Our main theorem are stated as follows:
Theorem 1.1.
Let the relaxation parameter $\tau$ satisfy
$\tau\leq\min\{\inf\limits_{z_{1}\in[v_{m},v_{-}]}\frac{\mu}{2|\sigma_{1}^{2}+p^{\prime}(z_{1})|},\inf\limits_{z_{2}\in[v_{m},v_{+}]}\frac{\mu}{2|\sigma_{2}^{2}+p^{\prime}(z_{2})|},1\}.$ (1.10)
For a given constant state $(v_{+},u_{+})\in\mathbb{R}_{+}\times\mathbb{R}$, there exist constants $\delta_{0},\varepsilon_{0}>0$ such that the following holds true.
For any constant states $(v_{-},u_{-})$ and $(v_{m},u_{m})$ satisfying (1.7) with
$|v_{+}-v_{m}|+|v_{m}-v_{-}|\leq\delta_{0}.$
Denote $(\widetilde{v}_{i},\widetilde{u}_{i},\widetilde{\Pi}_{i})(x-\sigma_{i}t)$ the i-viscous shock solution that are the traveling wave solutions for (1.3)-(1.4) for each $i=1,2$. Let $(v_{0},u_{0},\Pi_{0})$ be any initial data such that
±$\sum\limits_{\pm}\left(\|(v_{0}-v_{\pm},u_{0}-u_{\pm})\|_{L^{2}(\mathbb{R}_{\pm})}\right)+\|((v_{0})_{x},(u_{0})_{x})\|_{H^{1}(\mathbb{R})}+\sqrt{\tau}\|\Pi_{0}\|_{H^{2}(\mathbb{R})}<\varepsilon_{0},$ (1.11)
where $\mathbb{R}_{+}:=-\mathbb{R}_{-}=(0,+\infty)$. Then, the initial value problem (1.3)-(1.4) has a unique global-in-time solution $(v,u,\Pi)\in C^{1}((0,+\infty)\times\mathbb{R})$. Moreover, there exist an absolutely continuous shift $X_{i}(t)$ (defined in (2.10)) such that
$v(t,x)-\left(\widetilde{v}_{1}(x-\sigma_{1}t-X_{1}(t))+\widetilde{v}_{2}(x-\sigma_{2}t-X_{2}(t))-v_{m}\right)\in C(0,+\infty;H^{2}(\mathbb{R})),$ (1.12)
$u(t,x)-\left(\widetilde{u}_{1}(x-\sigma_{1}t-X_{1}(t))+\widetilde{u}_{2}(x-\sigma_{2}t-X_{2}(t))-u_{m}\right)\in C(0,+\infty;H^{2}(\mathbb{R})),$
$\Pi(t,x)-\left(\widetilde{\Pi}_{1}(x-\sigma_{1}t-X_{1}(t))+\widetilde{\Pi}_{2}(x-\sigma_{2}t-X_{2}(t))\right)\in C(0,+\infty;H^{2}(\mathbb{R})),$
and
$\sup\limits_{t\in[0,+\infty)}\left(\|v-\widetilde{v}\|^{2}_{H^{2}}+\|u-\widetilde{u}\|^{2}_{H^{2}}+\tau\|\Pi-\widetilde{\Pi}\|^{2}_{H^{2}}\right)$ (1.13)
$\qquad\qquad+\int_{0}^{+\infty}\left(\|\left((v-\widetilde{v})_{x},(u-\widetilde{u})_{x}\right)\|^{2}_{H^{1}}+\|\Pi-\widetilde{\Pi}\|^{2}_{H^{2}}\right)\mathrm{d}t$
$\leq C_{0}\left(\|v_{0}-\widetilde{v}_{0}(\cdot)\|^{2}_{H^{2}}+\|u_{0}-\widetilde{u}_{0}(\cdot)\|^{2}_{H^{2}}+\tau\|\Pi_{0}-\widetilde{\Pi}_{0}(\cdot)\|^{2}_{H^{2}}\right)+C_{0}\delta_{0},$
where $C_{0}$ is a universal constant independent of $\tau$. In addition, as $t\rightarrow+\infty$,
$\sup\limits_{x\in\mathbb{R}}$ $\Big{|}v(t,x)-\left(\widetilde{v}_{1}(x-\sigma_{1}t-X_{1}(t))+\widetilde{v}_{2}(x-\sigma_{2}t-X_{2}(t))-v_{m}\right),$ (1.14)
$u(t,x)-\left(\widetilde{u}_{1}(x-\sigma_{1}t-X_{1}(t))+\widetilde{u}_{2}(x-\sigma_{2}t-X_{2}(t))-u_{m}\right),$
$\sqrt{\tau}\left(\Pi(t,x)-\left(\widetilde{\Pi}_{1}(x-\sigma_{1}t-X_{1}(t))+\widetilde{\Pi}_{2}(x-\sigma_{2}t-X_{2}(t))\right)\right)\Big{|}\rightarrow 0,$
where
$\lim\limits_{t\rightarrow+\infty}|\dot{X}_{i}(t)|=0\qquad for\qquad i=1,2.$ (1.15)
In addition, the shifts are well-separated in the following sense:
$X_{1}(t)+\sigma_{1}t\leq\frac{\sigma_{1}t}{2}<0<\frac{\sigma_{2}t}{2}\leq X_{2}(t)+\sigma_{2}t,\quad t>0.$ (1.16)
Furthermore, based on the uniform estimates of error terms (1.13), we have the following convergence theorem.
Theorem 1.2.
Let $(v^{\tau},u^{\tau},\Pi^{\tau})$ be the global solutions obtained in Theorem 1.1. Then, there exists functions $(v^{0},u^{0})\in L^{\infty}\left((0,+\infty);H^{2}\right)$ and $\Pi^{0}\in L^{2}\left((0,+\infty);H^{2}\right)$, such that, as $\tau\rightarrow 0$
$(v^{\tau},u^{\tau})\rightharpoonup(v^{0},u^{0})\qquad weakly-*\quad in\quad L^{\infty}\left((0,+\infty);H^{2}\right),$
$\Pi^{\tau}\rightharpoonup\Pi^{0}\qquad weakly-\quad in\quad L^{2}\left((0,+\infty);H^{2}\right),$
where $(v^{0},u^{0})$ is the solution to the classical one-dimensional isentropic compressible Navier-Stokes equations (1.9), with initial value $(v_{0},u_{0})$. Moreover,
$\Pi^{0}=\mu\frac{(u^{0})_{x}}{v^{0}}.$
The structure of this paper is as follows. Some basic concept, including viscous shock wave and $a$-contraction with shifts theory are given in Section 2. In Section 3, we reformulate the original problem and present the a priori estimates (Proposition 3.1) which gives the proof of Theorem 1.1 immediately. In Section 4, we give a proof of Proposition 3.1. Finally, in Section 5, we prove that the solutions of relaxed system (1.3) converges globally in time to that of classical system (1.9).
Notations: $L^{p}(\mathbb{R})$ and $W^{s,p}(\mathbb{R})$ ($1\leq p\leq\infty$) denote the usual Lebesgue and Sobolev spaces over $\mathbb{R}$ with the norm $\|\cdot\|_{L^{p}}$ and $\|\cdot\|_{W^{s,p}}$, respectively. Note that, when $s=0$, $W^{0,p}=L^{p}$. For $p=2$, $W^{s,2}$ are abbreviated to $H^{s}$ as usual. Let $T$ and $B$ be a positive constant and a Banach space, respectively. $C^{k}(0,T;B)(k\geq 0)$ denotes the space of $B$-valued $k$-times continuously differentiable functions on $[0,T]$, and $L^{p}(0,T;B)$ denotes the space of $B$-valued $L^{p}$-functions on $[0,T]$. The corresponding space $B$-valued functions on $[0,\infty)$ are defined in an analogous manner.
2. Preliminaries
2.1. Traveling wave
In this part, we first show the existence of two traveling wave solutions for system (1.3). Let $\xi_{i}=x-\sigma_{i}t,i=1,2$, with $\sigma_{1}^{2}=\frac{p(v_{-})-p(v_{m})}{v_{m}-v_{-}}$ and $\sigma_{2}^{2}=\frac{p(v_{m})-p(v_{+})}{v_{+}-v_{m}}$ are the speed of 1-shock wave and 2-shock wave, respectively. Assume the functions $(\widetilde{u}_{i},\widetilde{v}_{i},\widetilde{\Pi}_{i})(\xi_{i})$ satisfy
$(\widetilde{u}_{1},\widetilde{v}_{1},\widetilde{\Pi}_{1})(\xi_{1})\rightarrow(v_{m},u_{m},0),\quad(\widetilde{u}_{2},\widetilde{v}_{2},\widetilde{\Pi}_{2})(\xi_{2})\rightarrow(v_{+},u_{+},0),\quad(\xi\rightarrow+\infty),$ (2.1)
$(\widetilde{u}_{1},\widetilde{v}_{1},\widetilde{\Pi}_{1})(\xi_{1})\rightarrow(v_{-},u_{-},0),\quad(\widetilde{u}_{2},\widetilde{v}_{2},\widetilde{\Pi}_{2})(\xi_{2})\rightarrow(v_{m},u_{m},0)\quad(\xi\rightarrow-\infty).$
Plugging the form $(\widetilde{u}_{i},\widetilde{v}_{i},\widetilde{\Pi}_{i})(\xi_{i})$ into system (1.3), we have the following ordinary differential equations
$\begin{cases}-\sigma_{i}(\widetilde{v}_{i})_{\xi_{i}}-(\widetilde{u}_{i})_{\xi_{i}}=0,\\ -\sigma_{i}(\widetilde{u}_{i})_{\xi_{i}}+p(\widetilde{v}_{i})_{\xi_{i}}=(\widetilde{\Pi}_{i})_{\xi_{i}},\\ -\sigma_{i}(\tau\widetilde{\Pi}_{i})_{\xi_{i}}+\widetilde{v}_{i}\widetilde{\Pi}_{i}=\mu(\widetilde{u}_{i})_{\xi_{i}},\end{cases}$ (2.2)
[…]
$\begin{cases}\sigma_{2}\widetilde{v}_{2}+\widetilde{u}_{2}=\sigma_{2}v_{m}+u_{m}=\sigma_{2}v_{+}+u_{+},\\ \widetilde{\Pi}_{2}=-\sigma_{2}(\widetilde{u}_{2}-u_{m})+(p(\widetilde{v}_{2})-p(v_{m})).\end{cases}$ (2.3)
Substituting (2.3) and (2.2)1italic-(2.2subscriptitalic-)$_{1}$ into (2.2)3italic-(2.2subscriptitalic-)$_{3}$, we derive that
$(\widetilde{v}_{2})_{\xi_{2}}=\frac{\widetilde{v}_{2}h_{2}(\widetilde{v}_{2})}{\mu\sigma_{2}+\tau\sigma_{2}h_{2}^{\prime}(\widetilde{v}_{2})},$ (2.4)
where $h_{2}(\widetilde{v}_{2})=\sigma_{2}^{2}(v_{m}-\widetilde{v}_{2})+(p(v_{m})-p(\widetilde{v}_{2}))$.
Similarly, for $i=1$, we have
$(\widetilde{v}_{1})_{\xi_{1}}=\frac{\widetilde{v}_{1}h_{1}(\widetilde{v}_{1})}{\mu\sigma_{1}+\tau\sigma_{1}h_{1}^{\prime}(\widetilde{v}_{1})},$ (2.5)
where $h_{1}(\widetilde{v}_{1})=\sigma_{1}^{2}(v_{-}-\widetilde{v}_{1})+(p(v_{-})-p(\widetilde{v}_{1}))$.
The following lemma show the existence and properties of solutions for (2.2).
Lemma 2.1.
Let (1.10) hold. For any states $(v_{-},u_{-})$, $(v_{m},u_{m})$, $(v_{+},u_{+})$, and $\sigma_{1}<0,\sigma_{2}>0$ satisfying R-H condition and Lax condition, there exists a positive constant $C$ independent of $\tau$ such that the following is true: the traveling wave solutions $(\widetilde{u}_{1},\widetilde{v}_{1},\widetilde{\Pi}_{1})(\xi_{1})$ connecting $(v_{-},u_{-},0)$ and $(v_{m},u_{m},0)$ and $(\widetilde{u}_{2},\widetilde{v}_{2},\widetilde{\Pi}_{2})(\xi_{2})$ connecting $(v_{m},u_{m},0)$ and $(v_{+},u_{+},0)$ exist uniquely and satisfy
$(\widetilde{v}_{1})_{\xi_{1}}<0,\quad(\widetilde{v}_{2})_{\xi_{2}}>0,\quad(\widetilde{v}_{i})_{\xi_{i}}\sim(\widetilde{u}_{i})_{\xi_{i}},$
and
$|\widetilde{v}_{i}(\xi_{i})-v_{m}|\leq C\delta_{i}e^{-C\delta_{i}|\xi_{i}|},|\widetilde{u}_{i}(\xi_{i})-u_{m}|\leq C\delta_{i}e^{-C\delta_{i}|\xi_{i}|},\quad\forall(-1)^{i}\xi_{i}<0,$
$|\widetilde{\Pi}_{i}|\leq C\delta_{i}^{2}e^{-C\delta_{i}|\xi_{i}|},\quad\quad|\partial_{\xi_{i}}(\widetilde{v}_{i},\widetilde{u}_{i})|\leq C\delta_{i}^{2}e^{-C\delta_{i}|\xi_{i}|},\,|\partial_{\xi_{i}}\widetilde{\Pi}_{i}|\leq C\delta_{i}|(\widetilde{v}_{i})_{\xi_{i}}|\quad\forall\xi_{i}\in\mathbb{R},$
$|\partial^{k}_{\xi_{i}}(\widetilde{v}_{i},\widetilde{u}_{i},\widetilde{\Pi}_{i})|\leq C\delta_{i}|\partial_{\xi_{i}}\widetilde{v}_{i}|,\qquad\forall\xi_{i}\in\mathbb{R},$
for $i=1,2$ and $k=2,3,4$, where $\delta_{i}$ denote the strength of the shock as $\delta_{1}:=|p(v_{-})-p(v_{m})|\sim|v_{-}-v_{m}|\sim|u_{-}-u_{m}|$ and $\delta_{2}:=|p(v_{m})-p(v_{+})|\sim|v_{m}-v_{+}|\sim|u_{m}-u_{+}|$.
Proof.
We only give a proof for case $i=1$. The case $i=2$ follows in a similar way. Firstly, we note that $\sigma^{2}_{1}=\frac{p(v_{-})-p(v_{m})}{v_{m}-v_{-}}$ and $h_{1}(\widetilde{v}_{1})=\sigma_{1}^{2}(v_{-}-\widetilde{v}_{1})+(p(v_{-})-p(\widetilde{v}_{1}))$, then we have
$h_{1}(\widetilde{v}_{1})=(v_{-}-\widetilde{v}_{1})\left(P(\widetilde{v}_{1})-P(v_{m})\right),$
where $P(\widetilde{v}_{1})=\frac{p(v_{-})-p(\widetilde{v}_{1})}{v_{-}-\widetilde{v}_{1}}$. The concavity of $p$ implies
$P^{\prime}(\widetilde{v}_{1})=\frac{-p^{\prime}(\widetilde{v}_{1})(v_{-}-\widetilde{v}_{1})+p(v_{-})-p(\widetilde{v}_{1})}{(v_{-}-\widetilde{v}_{1})^{2}}=\frac{p^{\prime\prime}(v_{\ast})}{2}>0,$
where $v_{\ast}$ is between $v_{-}$ and $v_{m}$. Therefore, we get $\widetilde{v}^{S}_{\xi}<0$.
In addition, using the entropy inequality: $-\sqrt{-p^{\prime}(v_{-})}>-\sqrt{-p^{\prime}(v_{m})}$, we get $v_{-}>v_{m}$. So, we derive that
$v_{m}<\widetilde{v}_{1}<v_{-}.$ (2.6)
Let $f(\widetilde{v}_{1})=\frac{\widetilde{v}_{1}h_{1}(\widetilde{v}_{1})}{\mu\sigma_{1}+\tau\sigma_{1}h^{\prime}_{1}(\widetilde{v}_{1})}$. Using (1.10), one has
$|f(v_{1})-f(v_{2})|$ $=\Big{|}\frac{v_{1}h(v_{1})}{\mu\sigma+\tau\sigma h^{\prime}(v_{1})}-\frac{v_{2}h(v_{2})}{\mu\sigma+\tau\sigma h^{\prime}(v_{2})}\Big{|}$
$\leq\Big{|}\frac{v_{1}h(v_{1})-v_{2}h(v_{2})}{\mu\sigma+\tau\sigma h^{\prime}(v_{1})}\Big{|}+\Big{|}v_{2}h(v_{2})\frac{\sigma\tau\left(h^{\prime}(v_{1})-h^{\prime}(v_{2})\right)}{(\mu\sigma+\tau\sigma h^{\prime}(v_{1}))(\mu\sigma+\tau\sigma h^{\prime}(v_{2}))}\Big{|}$
$\leq C\left(|v_{1}-v_{2}|+|h(v_{1})-h(v_{2})|+|h^{\prime}(v_{1})-h^{\prime}(v_{2})|\right)$
$\leq C|v_{1}-v_{2}|.$
So, $f(\widetilde{v}_{1})$ satisfies the usual Lipschitz condition. This together with (2.6) implies that there exists a unique global solution of (2.5)
Next, for $\xi_{1}<0$, using (2.5), we have
$(v_{-}-\widetilde{v}_{1})_{\xi_{1}}<\frac{2v_{-}p^{\prime\prime}(v_{\ast})(v_{-}-\widetilde{v}_{1})(\widetilde{v}_{1}-v_{m})}{-\mu\sigma_{1}}\leq C(v_{-}-\widetilde{v}_{1})(\widetilde{v}_{1}-v_{m})\leq C\delta_{1}(v_{-}-\widetilde{v}_{1}),$
which gives
$v_{-}-\widetilde{v}_{1}\leq C\delta_{1}e^{C\delta_{1}\xi_{1}}.$
On the other hand, if $\xi_{1}>0$, we have
$(v_{m}-\widetilde{v}_{1})_{\xi_{1}}>\frac{v_{m}p^{\prime\prime}(v_{\ast})(v_{-}-\widetilde{v}_{1})(\widetilde{v}_{1}-v_{m})}{2\mu\sigma_{1}}\geq C(v_{-}-\widetilde{v}_{1})(\widetilde{v}_{1}-v_{m})\geq C\delta_{1}(\widetilde{v}_{1}-v_{m}).$
Thus, it yields
$\widetilde{v}_{1}-v_{m}\leq C\delta_{1}e^{-C\delta_{1}\xi_{1}}.$
Combining the above estimates, one obtain
$(\widetilde{v}_{1})_{\xi_{1}}\leq C\delta_{1}^{2}e^{-C\delta_{1}|\xi_{1}|}.$
For high-order estimates, using (2.5) and (1.10), we have
$|(\widetilde{v}_{1})_{\xi_{1}\xi_{1}}|=\Big{|}\frac{(\widetilde{v}_{1})_{\xi_{1}}h((\widetilde{v}_{1})_{\xi_{1}})+\widetilde{v}_{1}h^{\prime}(\widetilde{v}_{1})(\widetilde{v}_{1})_{\xi_{1}}}{\mu\sigma_{1}+\tau\sigma_{1}h^{\prime}(\widetilde{v}_{1})}+\frac{\widetilde{v}_{1}h(\widetilde{v}_{1})\tau\sigma_{1}h^{\prime\prime}(\widetilde{v}_{1})(\widetilde{v}_{1})_{\xi_{1}}}{\left(\mu\sigma_{1}+\tau\sigma_{1}h^{\prime}(\widetilde{v}_{1})\right)^{2}}\Big{|}\leq C\delta_{1}|(\widetilde{v}_{1})_{\xi_{1}}|.$
Similarly, for $k=3,4$, we can get $|\partial_{\xi_{1}}^{k}\widetilde{v}_{1}|\leq C\delta_{1}|(\widetilde{v}_{1})_{\xi_{1}}|$.
Using (2.2)1italic-(2.2subscriptitalic-)$_{1}$, we get
$|\widetilde{u}_{1}-u_{-}|=|\sigma_{1}(\widetilde{v}_{1}-v_{-})|\leq C\delta_{1}e^{-C\delta_{1}|\xi_{1}|},\quad\xi_{1}<0,$
and
$|\widetilde{u}_{1}-u_{m}|=|\sigma_{1}(\widetilde{v}_{1}-v_{m})|\leq C\delta_{1}e^{-C\delta_{1}|\xi_{1}|},\quad\xi_{1}>0.$
Similarly, one has
$|(\widetilde{u}_{1})_{\xi_{1}}|=|\sigma_{1}(\widetilde{v}_{1})_{\xi_{1}}|\leq C\delta_{1}^{2}e^{-C\delta_{1}|\xi_{1}|},$
and
$|\partial_{\xi_{1}}^{k}\widetilde{u}_{1}|=|\sigma_{1}\partial_{\xi_{1}}^{k}\widetilde{v}_{1}|\leq C\delta_{1}|(\widetilde{v}_{1})_{\xi_{1}}|,$
where $k=2,3,4$.
[…]
$\widetilde{\Pi}_{1}=\frac{-\sigma_{1}^{3}\tau+\sigma_{1}\tau p^{\prime}(\widetilde{v}_{1})-\mu\sigma_{1}}{\widetilde{v}_{1}}(\widetilde{v}_{1})_{\xi},$ (2.7)
and thus
$|\widetilde{\Pi}_{1}|\leq C|(\widetilde{v}_{1})_{\xi}|.$ (2.8)
Taking the derivative of equation (2.7) with respect to $\xi_{1}$, we derive that
$(\widetilde{\Pi}_{1})_{\xi_{1}}=\frac{\sigma_{1}^{3}\tau+\sigma_{1}\tau p^{\prime}(\widetilde{v}_{1})-\mu\sigma_{1}}{\widetilde{v}_{1}}(\widetilde{v}_{1})_{\xi_{1}\xi_{1}}+\frac{\sigma_{1}\tau p^{\prime\prime}(\widetilde{v}_{1})(\widetilde{v}_{1})_{\xi_{1}}(\widetilde{v}_{1})-(\sigma_{1}^{3}\tau+\sigma_{1}\tau p^{\prime}(\widetilde{v}_{1})-\mu\sigma_{1})(\widetilde{v}_{1})_{\xi_{1}}}{(\widetilde{v}_{1})^{2}}(\widetilde{v}_{1})_{\xi_{1}}.$
Noting that $(\widetilde{v}_{1})_{\xi_{1}\xi_{1}}\leq C\delta_{1}(\widetilde{v}_{1})_{\xi_{1}}$, we have
$|(\widetilde{\Pi}_{1})_{\xi_{1}}|\leq C\delta_{1}(\widetilde{v}_{1})_{\xi_{1}}.$
Similarly, for $k=2,3,4$, we have $|\partial_{\xi}^{k}\widetilde{\Pi}_{1}|\leq C\delta_{1}(\widetilde{v}_{1})_{\xi_{1}}$. Thus, the proof of this lemma is finished. ∎
Now, we turn our attention to the composite waves formed by two viscous shocks. Initially, we note that for given end states $(v_{\pm},u_{\pm},0)\in\mathbb{R}^{+}\times\mathbb{R}\times\mathbb{R}$ as specified in (1.4), there exists a unique intermediate state $(v_{m},u_{m},0)$ such that $(v_{-},u_{-},0)$ is connected to $(v_{m},u_{m},0)$ by the 1-shock wave, and the 2-shock wave connects $(v_{m},u_{m},0)$ and $(v_{+},u_{+},0)$ for system (1.5)-(1.6), see [41]. The composite waves of the superposition of two viscous shocks is defined as follows:
$\left(\widetilde{v}_{1}(x-\sigma_{1}t)+\widetilde{v}_{2}(x-\sigma_{2}t)-v_{m},\widetilde{u}_{1}(x-\sigma_{1}t)+\widetilde{u}_{2}(x-\sigma_{2}t)-u_{m},\widetilde{\Pi}_{1}(x-\sigma_{1}t)+\widetilde{\Pi}_{2}(x-\sigma_{2}t)\right).$ (2.9)
2.2. Construction of shifts
We define the shifts $(X_{1}(t),X_{2}(t))$ as a solution to the following system of ODEs:
[…]
where $i=1,2$, $M$ is the specific constant chosen as $M:=\frac{5(\gamma+1)}{8\gamma p(v_{m})}(-p^{\prime}(v_{m}))^{\frac{3}{2}}$ and $f^{X_{i}(t)}$ denotes a function $f$ shifted by $X_{i}(t)$, that is $f^{X_{i}(t)}(x):=f(x-X_{i}(t))$. The shifted weight function $a$ is defined by
$a(t,x):=a_{1}(x-\sigma_{1}t-X_{1}(t))+a_{2}(x-\sigma_{2}t-X_{2}(t))-1,$ (2.11)
where $a_{1},a_{2}$ are two weight functions associated with 1-shock and 2-shock, respectively, defined by
$a_{i}(x-\sigma_{i}t)=1+\frac{\lambda_{i}(p(v_{m})-p(\widetilde{v}_{i}(x-\sigma_{i}t)))}{\delta_{i}},$
where $\lambda_{1},\lambda_{2}$ are small constants such that $\frac{\delta_{i}}{\lambda_{i}}<\delta_{0}$ and $\lambda_{i}<C\sqrt{\delta_{i}}<C\delta_{0}$ for $i=1,2$.
On the other hand, we known that $21<a<$ 2 and
$(a_{i})_{x}=-\frac{\lambda_{i}}{\delta_{i}}(p(\widetilde{v}_{i}))_{x}.$
Then, using Lemma 2.1, we derive that
$\sigma_{i}(a_{i})_{x}=-\sigma_{i}\frac{\lambda_{i}}{\delta_{i}}p^{\prime}(\widetilde{v}_{i})(\widetilde{v}_{i})_{x}>0$ (2.12)
and
$|(a_{i})_{x}|\leq C\frac{\lambda_{i}}{\delta_{i}}|\widetilde{v}_{i}|.$ (2.13)
In addition, based on the Cauchy-Lipschitz theorem, we have the following lemma, see [25, 10].
Lemma 2.2.
For any $c_{1},c_{2}>0$, there exists a constant $C>0$ such that the following is true. For any $T>0$, and any function $v\in L^{\infty}((0,T)\times\mathbb{R})$ veriying
$c_{1}<v(t,x)<c_{2},\quad\forall(t,x)\in[0,T]\times\mathbb{R},$ (2.14)
the ODE (2.10) has a unique absolutely continuous solution $(X_{1},X_{2})$ on $[0,T]$. Moreover,
[…]
We remark that the shift function above is slightly different with that in [10] due to the unavailability of BD entropy. Nonetheless, using similar proof as in [10], we can get that (2.10) has a unique absolutely continuous solution defined on any interval $[0,T]$ provided $v(t,x)$ is bounded both above and below for any $(t,x)\in[0,T]\times\mathbb{R}$. For completeness, we give a proof in Appendix A.
3. Local solution and a priori estimates
Firstly, by using the theory of symmetric hyperbolic system, we have the following local existence theory, see [17, 25, 33, 42].
Theorem 3.1.
Let $\underline{v}$ and $\underline{u}$ be smooth monotone functions such that
$\underline{v}(x)=v_{\pm}\quad\underline{u}(x)=u_{\pm}\quad for\quad\pm x\geq 1.$
For any constants $M_{0},M_{1},\kappa_{1},\kappa_{2},\kappa_{3},\kappa_{4}$ with $M_{1}>M_{0}>0$ and $\kappa_{1}>\kappa_{2}>\kappa_{3}>\kappa_{4}>0$, there exists a constant $T_{0}>0$ such that if
$\|v_{0}-\underline{v}\|_{H^{2}}+\|u_{0}-\underline{u}\|_{H^{2}}+\sqrt{\tau}\|\Pi_{0}\|_{H^{2}}\leq M_{0},$
$0<\kappa_{3}\leq v_{0}(x)\leq\kappa_{2},\qquad\forall x\in\mathbb{R},$
then the system (1.3) with initial condition (1.4) has a unique solution $(v,u,\Pi)$ on $[0,T_{0}]$ such that
$(v-\underline{v},u-\underline{u},\Pi)\in C([0,T_{0}];H^{2}),$
and
$\|v-\underline{v}\|_{L^{\infty}(0,T_{0};H^{2})}+\|u-\underline{u}\|_{L^{\infty}(0,T_{0};H^{2})}+\sqrt{\tau}\|\Pi\|_{L^{\infty}(0,T_{0};H^{2})}\leq M_{1}.$
Moreover:
$\kappa_{4}\leq v(t,x)\leq\kappa_{1},\qquad\forall(t,x)\in[0,T_{0}]\times\mathbb{R}.$
Next, let the shifted composite waves $(\widetilde{v},\widetilde{u},\widetilde{\Pi})(t,x)$ as follows:
$\left(\widetilde{v},\widetilde{u},\widetilde{\Pi}\right)(t,x):=\left(\widetilde{v}_{1}^{X_{1}}+\widetilde{v}_{2}^{X_{2}}-v_{m},\widetilde{u}_{1}^{X_{1}}+\widetilde{u}_{2}^{X_{2}}-u_{m},\widetilde{\Pi}_{1}^{X_{1}}+\widetilde{\Pi}_{2}^{X_{2}}\right),$ (3.1)
which satisfies the following shifted composite waves equations
$\begin{cases}\widetilde{v}_{t}+\sum\limits_{i=1}\limits^{2}\dot{X}_{i}(\widetilde{v}_{i})^{X_{i}}_{x}-\widetilde{u}_{x}=0,\\ \widetilde{u}_{t}+\sum\limits_{i=1}\limits^{2}\dot{X}_{i}(\widetilde{u}_{i})^{X_{i}}_{x}+p(\widetilde{v})_{x}=\widetilde{\Pi}_{x}+F_{1},\\ \tau\widetilde{\Pi}_{t}+\tau\sum\limits_{i=1}\limits^{2}\dot{X}_{i}(\widetilde{\Pi}_{i})^{X_{i}}_{x}+\widetilde{v}\widetilde{\Pi}=\mu\widetilde{u}_{x}+F_{2},\end{cases}$ (3.2)
where $F_{1},F_{2}$ defined as
$F_{1}:=p(\widetilde{v})_{x}-p(\widetilde{v}_{1})^{X_{1}}_{x}-p(\widetilde{v}_{2})^{X_{2}}_{x},\quad F_{2}:=\left(\widetilde{v}^{X_{2}}_{2}-v_{m}\right)\widetilde{\Pi}_{1}^{X_{1}}+\left(\widetilde{v}^{X_{1}}_{1}-v_{m}\right)\widetilde{\Pi}_{2}^{X_{2}}.$ (3.3)
Next, we will focus on the time-asymptotic stability of the solutions to (1.3) with initial data (1.4) to the the shifted composite waves for system (3.2). Denote $\left(\widetilde{v}_{0}(\cdot),\widetilde{u}_{0}(\cdot),\widetilde{\Pi}_{0}(\cdot)\right)=\left(\widetilde{v},\widetilde{u},\widetilde{\Pi}\right)(t,\cdot)\Big{|}_{t=0}$. Then, one can easily get
±$\sum\limits_{\pm}\left(\|(\widetilde{v}_{0}-v_{\pm},\widetilde{u}_{0}-u_{\pm},\sqrt{\tau}\widetilde{\Pi}_{0})\|_{H^{2}(\mathbb{R}_{\pm})}\right)\leq C(\delta_{1}+\delta_{2}).$ (3.4)
The following proposition gives the a priori estimates of the error term: $(v-\widetilde{v},u-\widetilde{u},\Pi-\widetilde{\Pi})$.
Proposition 3.1.
Let $(v,u,\Pi)$ be local solutions given by Theorem 3.1 on $[0,T]$ for some $T>0$ and $(\widetilde{v},\widetilde{u},\widetilde{\Pi})$ be defined in (3.1). Then there exist positive constants $\delta_{0},\varepsilon_{1}$ such that if two independent shock waves strength satisfy $\delta_{1},\delta_{2}<\delta_{0}$ and
$\sup_{0\leq t\leq T}\|(v-\widetilde{v},u-\widetilde{u},\sqrt{\tau}(\Pi-\widetilde{\Pi}))\|_{H^{2}}\leq\varepsilon_{1},$ (3.5)
then the following estimates hold
$\sup\limits_{t\in[0,T]}\left(\|v-\widetilde{v}\|^{2}_{H^{2}}+\|u-\widetilde{u}\|^{2}_{H^{2}}+\tau\|\Pi-\widetilde{\Pi}\|^{2}_{H^{2}}\right)+\int_{0}^{T}\sum\limits_{i=1}\limits^{2}\delta_{i}|\dot{X}_{i}|^{2}\mathrm{d}t$ (3.6)
$\qquad\qquad+\int_{0}^{T}\left(G^{s}(U)+\|\left((v-\widetilde{v})_{x},(u-\widetilde{u})_{x}\right)\|^{2}_{H^{1}}+\|\Pi-\widetilde{\Pi}\|^{2}_{H^{2}}\right)\mathrm{d}t$
$\leq C_{0}\left(\|v_{0}-\widetilde{v}_{0}(\cdot)\|^{2}_{H^{2}}+\|u_{0}-\widetilde{u}_{0}(\cdot)\|^{2}_{H^{2}}+\tau\|\Pi_{0}-\widetilde{\Pi}_{0}(\cdot)\|^{2}_{H^{2}}\right)+C_{0}\delta_{0},$
where $C_{0}$ independent of $T$ and $\tau$ and
$G^{s}(U):=\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||v-\widetilde{v}|^{2}\mathrm{d}x$ (3.7)
In addition, by (2.10),
$|\dot{X}_{1}(t)|+|\dot{X}_{2}(t)|\leq C_{0}\|(v-\widetilde{v})(t,\cdot)\|_{L^{\infty}},\qquad\forall t\leq T.$ (3.8)
We will give the proof of Proposition 3.1 in Section 4.
Now, by using Theorem 3.1 and Proposition 3.1, we are able to prove Theorem 1.1.
Proof of Theorem 1.1: Firstly, by definition of $\underline{v},\underline{u}$, for some $C_{\ast}>0$, we have
±$\sum\limits_{\pm}\left(\|(\underline{v}-v_{\pm},\underline{u}-u_{\pm})\|_{L^{2}(\mathbb{R}_{\pm})}\right)+\|((\underline{v})_{x},(\underline{u})_{x})\|_{H^{1}}<C_{\ast}(\delta_{1}+\delta_{2}).$ (3.9)
Then, using (3.4), we can derive that for some $C_{\ast\ast}>0$,
$\|(\underline{v}-\widetilde{v}_{0},\underline{u}-\widetilde{u}_{0})\|_{H^{2}}+\sqrt{\tau}\|\widetilde{\Pi}_{0}\|_{H^{2}}$ (3.10)
±$\leq\sum\limits_{\pm}\left(\|(\underline{v}-v_{\pm},\underline{u}-u_{\pm})\|_{H^{2}(\mathbb{R}_{\pm})}\right)+\sum\limits_{\pm}\left(\|(\widetilde{v}_{0}-v_{\pm},\widetilde{u}_{0}-u_{\pm},\sqrt{\tau}\widetilde{\Pi}_{0})\|_{H^{2}(\mathbb{R}_{\pm})}\right)$
$<C_{\ast\ast}(\sqrt{\delta_{1}}+\sqrt{\delta_{2}}).$
Noting that $\delta_{1},\delta_{2}\in(0,\delta_{0})$ and using the smallness of $\delta_{0}$, we get
$\frac{\frac{\varepsilon_{1}}{2}-C_{0}\delta_{0}}{C_{0}+1}-C_{\ast\ast}(\sqrt{\delta_{1}}+\sqrt{\delta_{2}})-C_{\ast}(\delta_{1}+\delta_{2})>0.$ (3.11)
Then, based on the above positive constants, we define $\varepsilon_{0}$:
$\varepsilon_{0}:=\varepsilon_{\ast}-C_{\ast}(\delta_{1}+\delta_{2}),\qquad\varepsilon_{\ast}:=\frac{\frac{\varepsilon_{1}}{2}-C_{0}\delta_{0}}{C_{0}+1}-C_{\ast\ast}(\sqrt{\delta_{1}}+\sqrt{\delta_{2}}),$
where we can choose $\varepsilon_{0}$ independent of $\delta_{1},\delta_{2}$. For example, we can take $\varepsilon_{0}=\frac{\varepsilon_{1}}{4(C_{0}+1)}$.
Now, by using the assumption (1.11) in Theorem 1.1 and (3.9), we can get
$\|(v_{0}-\underline{v},u_{0}-\underline{u})\|_{H^{2}}+\sqrt{\tau}\|\Pi_{0}\|_{H^{2}}$ (3.12)
±$\leq\sum\limits_{\pm}\left(\|(v_{0}-v_{\pm},u_{0}-u_{\pm},\sqrt{\tau}\Pi_{0})\|_{H^{2}(\mathbb{R}_{\pm})}\right)+\sum\limits_{\pm}\left(\|(\underline{v}-v_{\pm},\underline{u}-u_{\pm})\|_{H^{2}(\mathbb{R}_{\pm})}\right)$
$<\varepsilon_{0}+C_{\ast}(\delta_{1}+\delta_{2})=\varepsilon_{\ast}.$
Especially, by applying Sobolev’s embedding theorem, we obtain
$\|v_{0}-\underline{v}\|_{L^{\infty}}\leq C\varepsilon_{\ast},$
and subsequently, using the smallness of $\varepsilon_{\ast}$, we can derive that
$\frac{v_{-}}{2}<v_{0}(x)<2v_{+},\qquad\forall x\in\mathbb{R}.$ (3.13)
Applying the smallness of $\delta_{1},\delta_{2}$, we have
$0<\varepsilon_{\ast}<\frac{\varepsilon_{1}}{2}.$ (3.14)
Therefore, according Theorem 3.1 with (3.12), (3.13) and (3.14), we obtain that the system (1.3) with initial condition (1.4) has a unique solution $(v,u,\Pi)$ on $[0,T_{0}]$ such that
$(v-\underline{v},u-\underline{u},\Pi)\in C([0,T_{0}];H^{2}),$ (3.15)
$\|v-\underline{v}\|_{L^{\infty}(0,T_{0};H^{2})}+\|u-\underline{u}\|_{L^{\infty}(0,T_{0};H^{2})}+\sqrt{\tau}\|\Pi\|_{L^{\infty}(0,T_{0};H^{2})}\leq\frac{\varepsilon_{1}}{2}$ (3.16)
and
$\frac{v_{-}}{3}\leq v(t,x)\leq 3v_{+},\qquad\forall(t,x)\in[0,T_{0}]\times\mathbb{R}.$ (3.17)
In addition, using the definition of $(\widetilde{v},\widetilde{u},\widetilde{\Pi})$ and Lemma 2.1, we have
$\int_{0}^{\infty}|\widetilde{v}_{2}(x-\sigma_{2}t-X_{2}(t))-v_{+}|^{2}\mathrm{d}x$
$\leq\int_{0}^{\infty}|\widetilde{v}_{2}(x-\sigma_{2}t)-v_{+}|^{2}\mathrm{d}x+\int_{-X_{2}(t)}^{0}|\widetilde{v}_{2}(x-\sigma_{2}t)-v_{+}|^{2}\mathrm{d}x$
$\leq C\delta_{2}(1+|X_{2}(t)|).$
Similarly, we can get
$\|\widetilde{v}_{2}-v_{m}\|_{L^{2}(\mathbb{R}_{-})}^{2}\leq C\delta_{2}(1+|X_{2}(t)|),$
and
$\|\widetilde{v}_{1}-v_{-}\|_{L^{2}(\mathbb{R}_{-})}^{2}\leq C\delta_{1}(1+|X_{1}(t)|),\,\|\widetilde{v}_{1}-v_{m}\|_{L^{2}(\mathbb{R}_{+})}^{2}\leq C\delta_{1}(1+|X_{1}(t)|).$
Hence, combining the above estimates and using Lemma 2.1, one get
$\|(\underline{v}-\widetilde{v},\underline{u}-\widetilde{u})\|_{H^{2}}+\sqrt{\tau}\|\widetilde{\Pi}\|_{H^{2}}$
±$\leq\sum\limits_{\pm}\left(\|(\underline{v}-v_{\pm},\underline{u}-u_{\pm})\|_{H^{2}(\mathbb{R}_{\pm})}\right)+\sum\limits_{\pm}\left(\|(\widetilde{v}-v_{\pm},\widetilde{u}-u_{\pm},\sqrt{\tau}\widetilde{\Pi})\|_{H^{2}(\mathbb{R}_{\pm})}\right)$
$<C\sqrt{\delta_{1}}(1+\sqrt{|X_{1}(t)|})+C\sqrt{\delta_{2}}(1+\sqrt{|X_{2}(t)|}).$
Then, using Lemma 2.2, we can get
$\|(\underline{v}-\widetilde{v},\underline{u}-\widetilde{u})\|_{H^{2}}+\sqrt{\tau}\|\widetilde{\Pi}\|_{H^{2}}<C\sqrt{\delta_{0}}(1+\sqrt{t}).$ (3.18)
Next, choosing $0<T_{1}<T_{0}$ small enough such that $C\sqrt{\delta_{0}}(1+\sqrt{t})<\frac{\varepsilon_{1}}{2}$, we have
$\|(\underline{v}-\widetilde{v},\underline{u}-\widetilde{u})\|_{L^{\infty}(0,T_{1};H^{2})}+\sqrt{\tau}\|\widetilde{\Pi}\|_{L^{\infty}(0,T_{1};H^{2})}<\frac{\varepsilon_{1}}{2}.$ (3.19)
Combining (3.16) and (3.19), we obtain that
$\|(v-\widetilde{v},u-\widetilde{u})\|_{L^{\infty}(0,T_{1};H^{2})}+\sqrt{\tau}\|\Pi-\widetilde{\Pi}\|_{L^{\infty}(0,T_{1};H^{2})}<\varepsilon_{1}.$ (3.20)
Additionally, since the shifts $X_{i}(t)$ are Lipschitz continuous and using (3.20), we can get
$(v-\widetilde{v},u-\widetilde{u},\Pi-\widetilde{\Pi})\in C([0,T_{1}];H^{2}),$ (3.21)
Now, we consider the maximal existence time defined as follows
$T_{m}:=\sup\left\{t>0\Big{|}\sup\limits_{[0,t]}\left(\|(v-\widetilde{v},u-\widetilde{u})\|_{H^{2}}+\sqrt{\tau}\|\Pi-\widetilde{\Pi}\|_{H^{2}}\right)\leq\varepsilon_{1}\right\}.$
If $T_{m}<\infty$, according the continuation argument, we can obtain that
$\sup\limits_{[0,T_{m}]}\left(\|(v-\widetilde{v},u-\widetilde{u})\|_{H^{2}}+\sqrt{\tau}\|\Pi-\widetilde{\Pi}\|_{H^{2}}\right)=\varepsilon_{1}.$ (3.22)
However, by (3.10), (3.11) and (3.12), we can get
$\|(v_{0}-\widetilde{v}_{0},u_{0}-\widetilde{u}_{0})\|_{H^{2}}+\sqrt{\tau}\|\Pi_{0}-\widetilde{\Pi}_{0}\|_{H^{2}}\leq\frac{\frac{\varepsilon_{1}}{2}-C_{0}\delta_{0}}{C_{0}+1}.$ (3.23)
Next, based on the Proposition 3.1 and smallness of $\delta_{0}$, it can be deduced that
$\sup\limits_{[0,T_{m}]}\left(\|(v-\widetilde{v},u-\widetilde{u})\|_{H^{2}}+\sqrt{\tau}\|\Pi-\widetilde{\Pi}\|_{H^{2}}\right)$ $\leq C_{0}\frac{\frac{\varepsilon_{1}}{2}-C_{0}\delta_{0}}{C_{0}+1}+C_{0}\delta_{0}$ (3.24)
$\leq\frac{\varepsilon_{1}}{2},$
which contradicts (3.22).
Hence, we conclude that $T_{m}=\infty$ and together with Proposition 3.1, we can derive that
$\sup\limits_{t\in[0,\infty)}\left(\|v-\widetilde{v}\|^{2}_{H^{2}}+\|u-\widetilde{u}\|^{2}_{H^{2}}+\tau\|\Pi-\widetilde{\Pi}\|^{2}_{H^{2}}\right)+\int_{0}^{\infty}\sum\limits_{i=1}\limits^{2}\delta_{i}|\dot{X}_{i}|^{2}\mathrm{d}t$ (3.25)
$\qquad\qquad+\int_{0}^{\infty}\left(G^{s}(U)+\|\left((v-\widetilde{v})_{x},(u-\widetilde{u})_{x}\right)\|^{2}_{H^{1}}+\|\Pi-\widetilde{\Pi}\|^{2}_{H^{2}}\right)\mathrm{d}t$
$\leq C_{0}\left(\|v_{0}-\widetilde{v}_{0}(\cdot)\|^{2}_{H^{2}}+\|u_{0}-\widetilde{u}_{0}(\cdot)\|^{2}_{H^{2}}+\tau\|\Pi_{0}-\widetilde{\Pi}_{0}(\cdot)\|^{2}_{H^{2}}\right)+C_{0}\delta_{0}.$
In addition, using system (4.1) below and (3.25), we can get
$\int_{0}^{\infty}\left\|\partial_{x}(v-\widetilde{v},u-\widetilde{u},\Pi-\widetilde{\Pi})\right\|_{L^{2}}^{2}\mathrm{d}t\leq C$
and
$\int_{0}^{\infty}\Big{|}\frac{\mathrm{d}}{\mathrm{d}t}\left\|\partial_{x}(v-\widetilde{v},u-\widetilde{u},\sqrt{\tau}(\Pi-\widetilde{\Pi}))\right\|_{L^{2}}^{2}\Big{|}\mathrm{d}t\leq C.$
Thus, combining the interpolation inequality, we get
$\lim\limits_{t\rightarrow\infty}\left\|(v-\widetilde{v},u-\widetilde{u},\sqrt{\tau}(\Pi-\widetilde{\Pi}))\right\|_{L^{\infty}}=0.$
Furthermore, using the definition of shifts $X_{1}(t),X_{2}(t)$ (see (2.10)), we have
$|\dot{X}_{1}(t)|+|\dot{X}_{2}(t)|\leq C\left\|v-\widetilde{v}\right\|_{L^{\infty}}\rightarrow 0\quad as\quad t\rightarrow\infty.$ (3.26)
On the other hand, using (3.25) and Sobolev’s embedding theorem, we have
$\|v-\widetilde{v}\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}\leq C_{0}(\varepsilon_{1}+\delta_{0}).$
Thus, using (3.26) and the smallness of $\varepsilon_{1}$ and $\delta_{0}$, we get
$X_{1}(t)\leq-\frac{\sigma_{1}}{2}t,\qquad X_{2}(t)\geq-\frac{\sigma_{2}}{2}t,\qquad t>0,$
or equivalently,
$X_{1}(t)+\sigma_{1}t\leq\frac{\sigma_{1}}{2}t,\qquad X_{2}(t)+\sigma_{2}t\geq\frac{\sigma_{2}}{2}t,\qquad t>0.$
Therefore, the proof of Theorem 1.1 is finished.
4. Proof of the Proposition 3.1
In this section, we establish the a priori estimates of local solutions and thus give a proof of Proposition 3.1. The constant $C$ (may different in different place) denotes a universal constant independent of $\tau$ and $T$.
Using (1.3) and (3.2), we derive the equations for error term $(v-\widetilde{v},u-\widetilde{u},\Pi-\widetilde{\Pi})(t,x)$ as follows:
$\begin{cases}(v-\widetilde{v})_{t}-\sum\limits_{{i}={1}}^{2}\dot{X}_{i}\cdot(\widetilde{v}_{i})^{X_{i}}_{x}-(u-\widetilde{u})_{x}=0,\\ (u-\widetilde{u})_{t}-\sum\limits_{{i}={1}}^{2}\dot{X}_{i}\cdot(\widetilde{u}_{i})^{X_{i}}_{x}+\left(p(v)-p(\widetilde{v})\right)_{x}=(\Pi-\widetilde{\Pi})_{x}-F_{1},\\ \tau(\Pi-\widetilde{\Pi})_{t}-\tau\sum\limits_{{i}={1}}^{2}\dot{X}_{i}\cdot(\widetilde{\Pi}_{i})^{X_{i}}_{x}+v\Pi-\widetilde{v}\widetilde{\Pi}=\mu(u-\widetilde{u})_{x}-F_{2},\end{cases}$ (4.1)
where $F_{1},F_{2}$ are defined in (3.3).
To prove Proposition 3.1, we need to do the lower-order, higher-order and dissipation estimates, respectively.
Firstly, let $U(t,x):=(v,u,\Pi)^{T}(t,x)$ and $\widetilde{U}(t,x):=(\widetilde{v},\widetilde{u},\widetilde{\Pi})^{T}(t,x)$ are the solution of system (1.3) and (3.2), respectively. Then, we define a relative entropy quantity as
$\eta(U|\widetilde{U})=\frac{|u-\widetilde{u}|^{2}}{2}+H(v|\widetilde{v})+\frac{\tau|\Pi-\widetilde{\Pi}|^{2}}{2\mu},$ (4.2)
where $H(v|\widetilde{v})=H(v)-H(\widetilde{v})-H^{\prime}(\widetilde{v})(v-\widetilde{v})$ and potential energy $H(v):=v^{-\gamma+1}/(\gamma-1)$. In addition, for the pressure function $p(v)=v^{-\gamma}$, denote by $p(v|\tilde{v})=p(v)-p(\widetilde{v})-p^{\prime}(\widetilde{v})(v-\widetilde{v})$. Then, we have the following Lemma, see [25, 10].
Lemma 4.1.
For given constants $\gamma>1$, and $v_{-}>0$, there exist constants $C,\delta_{\ast}>0$, such that the following holds true.
  • 1)
    For any $v,w$ such that $0<w<2v_{-},0<v<3v_{-}$,
    $|v-w|^{2}\leq CH(v|w),$
    $|v-w|^{2}\leq Cp(v|w).$
  • 2)
    For any $v,w>v_{-}/2$,
    $|p(v)-p(w)|\leq C|v-w|.$
  • 3)
    For any $0<\delta<\delta_{\ast}$, and for any $(v,w)\in\mathbb{R}^{2}_{+}$ satisfying $|p(v)-p(w)|<\delta$, and $|p(w)-p(v_{-})|<\delta$, the following holds true:
    $p(v|w)\leq\left(\frac{\gamma+1}{2\gamma}\frac{1}{p(w)}+C\delta\right)|p(v)-p(w)|^{2},$
    $H(v|w)\geq\frac{p(w)^{-\frac{1}{\gamma}-1}}{2\gamma}|p(v)-p(w)|^{2}-\frac{1+\gamma}{3\gamma^{2}}p(w)^{-\frac{1}{\gamma}-2}(p(v)-p(w))^{3},$
    $H(v|w)\leq\left(\frac{p(w)^{-\frac{1}{\gamma}-1}}{2\gamma}+C\delta\right)|p(v)-p(w)|^{2}.$
4.1. $L^{2}$ energy estimates
In the following lemmas, we get the $L^{2}$ estimates.
Lemma 4.2.
Under the hypotheses of Proposition 3.1, there exist $C,C_{2}>0$ (independent of $\tau$ and $T$) such that for all $t\in(0,T]$,
$\|v-\widetilde{v}\|^{2}_{L^{2}}+\|u-\widetilde{u}\|^{2}_{L^{2}}+\tau\|\Pi-\widetilde{\Pi}\|^{2}_{L^{2}}+\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}|\dot{X}_{i}|^{2}\mathrm{d}t+\int_{0}^{t}\left(G^{s}(U)+G(U)\right)\mathrm{d}t$ (4.3)
$\leq C\left(\|v_{0}-\widetilde{v}_{0}(\cdot)\|^{2}_{L^{2}}+\|u_{0}-\widetilde{u}_{0}(\cdot)\|^{2}_{L^{2}}+\tau\|\Pi_{0}-\widetilde{\Pi}_{0}(\cdot)\|^{2}_{L^{2}}\right)+C_{2}\int_{0}^{t}\int_{\mathbb{R}}\left|\partial_{x}(v-\widetilde{v})\right|^{2}\mathrm{d}x\mathrm{d}t+C\delta_{0},$
where
$G(U):=\int_{\mathbb{R}}\frac{v}{\mu}|\Pi-\widetilde{\Pi}|^{2}\mathrm{d}x,$ (4.4)
and $G^{s}(U)$ is given in (3.7).
Before we prove Lemma 4.2, we need some auxiliary lemmas. Firstly, using Lemma 2.1, we have the following estimates.
Next, using the a priori assumption (3.5), Sobolev’s embedding theorem and (2.10), we get that
$|\dot{X}_{i}(t)|\leq C\|v-\widetilde{v}\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}\leq C\varepsilon_{1}.$
Thus, for sufficiently small $\varepsilon_{1}$, we can derive that
$X_{1}(t)+\sigma_{1}t\leq\frac{\sigma_{1}}{2}t,\qquad X_{2}(t)+\sigma_{2}t\geq\frac{\sigma_{2}}{2}t,\qquad t>0.$ (4.5)
Lemma 4.3.
Assume (4.5). Given $v_{+}>0$, there exist positive constants $\delta_{0},C$ such that for any $\delta_{1},\delta_{2}\in(0,\delta_{0})$, the following estimates hold. For each $i=1,2$,
$|(\widetilde{v}_{i})_{x}^{X_{i}}||\widetilde{v}-\widetilde{v}_{i}^{X_{i}}|\leq C\delta_{i}\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t},\quad t>0,\quad x\in\mathbb{R},$
$\int_{\mathbb{R}}|(\widetilde{v}_{i})_{x}^{X_{i}}||\widetilde{v}-\widetilde{v}_{i}^{X_{i}}|\mathrm{d}x\leq C\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t},\quad t>0,$
$\int_{\mathbb{R}}|(\widetilde{v}_{1})_{x}^{X_{1}}||(\widetilde{v}_{2})_{x}^{X_{2}}|\mathrm{d}x\leq C\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t},\quad t>0.$
The proof of the above lemma which is based on Lemma 2.1 is given in [10]. For completeness, we give a proof of this lemma in the Appendix B.
Then, following the Lemma 4.3 in [8] (see also Lemma 4.3 in [25], Lemma 5.2 in [22] and Lemma 4 in [43]), we get the estimates of the relative entropy weighted by $a(t,x)$ with shift $X_{i}$.
Lemma 4.4.
Let $a(t,x)$ be the weight function defined in (2.11) and $X_{1},X_{2}$ be any Lipschitz continuous function. Let $U$ and $\widetilde{U}$ are the solution of system (1.3) and (3.2). Then, we have
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}}a(t,x)\eta(U(t,x)|\widetilde{U}(t,x))\mathrm{d}x=\sum\limits_{i=1}\limits^{2}\dot{X}_{i}(t)Y_{i}(U)+J^{bad}(U)-J^{good}(U),$ (4.6)
where
$Y_{i}(U):=$ $-\int_{\mathbb{R}}(a_{i})_{x}^{X_{i}}\eta(U|\widetilde{U})\mathrm{d}x+\int_{\mathbb{R}}a(\widetilde{u}_{i})^{X_{i}}_{x}(u-\widetilde{u})\mathrm{d}x-\int_{\mathbb{R}}ap^{\prime}(\widetilde{v})(\widetilde{v}_{i})^{X_{i}}_{x}(v-\widetilde{v})\mathrm{d}x$
$+\int_{\mathbb{R}}a\frac{\tau}{\mu}(\widetilde{\Pi}_{i})^{X_{i}}_{x}(\Pi-\widetilde{\Pi})\mathrm{d}x,$
$J^{bad}(U):=$ $\sum\limits_{i=1}\limits^{2}\left(\frac{1}{2\sigma_{i}}\int_{\mathbb{R}}(a_{i})_{x}^{X_{i}}(p(v)-p(\widetilde{v}))^{2}\mathrm{d}x+\sigma_{i}\int_{\mathbb{R}}a(\widetilde{v}_{i})^{X_{i}}_{x}p(v|\widetilde{v})\mathrm{d}x\right)$
$+\sum\limits_{i=1}\limits^{2}\left(\frac{1}{2\sigma_{i}}\int_{\mathbb{R}}(a_{i})_{x}^{X_{i}}(\Pi-\widetilde{\Pi})^{2}\mathrm{d}x-\frac{1}{2\sigma_{i}}\int_{\mathbb{R}}(a_{i})^{X_{i}}_{x}(p(v)-p(\widetilde{v}))(\Pi-\widetilde{\Pi})\mathrm{d}x\right)$
$-\int_{\mathbb{R}}a\frac{\widetilde{\Pi}}{\mu}(\Pi-\widetilde{\Pi})(v-\widetilde{v})\mathrm{d}x-\int_{\mathbb{R}}a(u-\widetilde{u})F_{1}\mathrm{d}x-\int_{\mathbb{R}}a\frac{(\Pi-\widetilde{\Pi})}{\mu}F_{2}\mathrm{d}x,$
$J^{good}(U):=$ $\sum\limits_{i=1}\limits^{2}\left(\frac{\sigma_{i}}{2}\int_{\mathbb{R}}(a_{i})_{x}^{X_{i}}\left(u-\widetilde{u}-\frac{p(v)-p(\widetilde{v})}{\sigma_{i}}+\frac{\Pi-\widetilde{\Pi}}{\sigma_{i}}\right)^{2}\mathrm{d}x+\sigma_{i}\int_{\mathbb{R}}(a_{i})^{X_{i}}_{x}H(v|\widetilde{v})\mathrm{d}x\right)$
$+\sum\limits_{i=1}\limits^{2}\left(\sigma_{i}\int_{\mathbb{R}}(a_{i})^{X_{i}}_{x}\frac{\tau}{2\mu}(\Pi-\widetilde{\Pi})^{2}\mathrm{d}x\right)+\int_{\mathbb{R}}a\frac{v}{\mu}(\Pi-\widetilde{\Pi})^{2}\mathrm{d}x.$
Next, for the terms $J^{bad}(U)$ and $J^{good}(U)$, we use the following notations:
$J^{bad}(U):=\sum\limits_{j=1}\limits^{7}B_{j}(U),\quad J^{good}(U):=G(U)+\sum\limits_{j=1}\limits^{3}G_{j}(U),$
where
$B_{1}(U):=\sum\limits_{i=1}\limits^{2}\left(\frac{1}{2\sigma_{i}}\int_{\mathbb{R}}(a_{i})_{x}^{X_{i}}(p(v)-p(\widetilde{v}))^{2}\mathrm{d}x\right),\quad B_{2}(U):=\sum\limits_{i=1}\limits^{2}\left(\sigma_{i}\int_{\mathbb{R}}a(\widetilde{v}_{i})^{X_{i}}_{x}p(v|\widetilde{v})\mathrm{d}x\right),$
$B_{3}(U):=\sum\limits_{i=1}\limits^{2}\left(\frac{1}{2\sigma_{i}}\int_{\mathbb{R}}(a_{i})_{x}^{X_{i}}(\Pi-\widetilde{\Pi})^{2}\mathrm{d}x\right),B_{4}(U):=-\sum\limits_{i=1}\limits^{2}\left(\frac{1}{2\sigma_{i}}\int_{\mathbb{R}}(a_{i})^{X_{i}}_{x}(p(v)-p(\widetilde{v}))(\Pi-\widetilde{\Pi})\mathrm{d}x\right),$
$B_{5}(U):=-\int_{\mathbb{R}}a\frac{\widetilde{\Pi}}{\mu}(\Pi-\widetilde{\Pi})(v-\widetilde{v})\mathrm{d}x,B_{6}(U):=-\int_{\mathbb{R}}a(u-\widetilde{u})F_{1}\mathrm{d}x,B_{7}(U):=-\int_{\mathbb{R}}a\frac{(\Pi-\widetilde{\Pi})}{\mu}F_{2}\mathrm{d}x,$
and
$G(U):=\int_{\mathbb{R}}a\frac{v}{\mu}(\Pi-\widetilde{\Pi})^{2}\mathrm{d}x,G_{1}(U):=\sum\limits_{i=1}\limits^{2}\left(\frac{\sigma_{i}}{2}\int_{\mathbb{R}}(a_{i})_{x}^{X_{i}}\left(u-\widetilde{u}-\frac{p(v)-p(\widetilde{v})}{\sigma_{i}}+\frac{\Pi-\widetilde{\Pi}}{\sigma_{i}}\right)^{2}\mathrm{d}x\right),$
$G_{2}(U):=\sum\limits_{i=1}\limits^{2}\sigma_{i}\int_{\mathbb{R}}(a_{i})^{X_{i}}_{x}H(v|\widetilde{v})\mathrm{d}x,G_{3}(U):=\sum\limits_{i=1}\limits^{2}\left(\sigma_{i}\int_{\mathbb{R}}(a_{i})^{X_{i}}_{x}\frac{\tau}{2\mu}(\Pi-\widetilde{\Pi})^{2}\mathrm{d}x\right).$
For each $Y_{i}(U)$, we have from (4.2) that
$Y_{i}(U):=$ $-\int_{\mathbb{R}}(a_{i})_{x}^{X_{i}}\left(\frac{|u-\widetilde{u}|^{2}}{2}+H(v|\widetilde{v})+\frac{\tau|\Pi-\widetilde{\Pi}|^{2}}{2\mu}\right)\mathrm{d}x+\int_{\mathbb{R}}a(\widetilde{u}_{i})^{X_{i}}_{x}(u-\widetilde{u})\mathrm{d}x$
$-\int_{\mathbb{R}}ap^{\prime}(\widetilde{v})(\widetilde{v}_{i})^{X_{i}}_{x}(v-\widetilde{v})\mathrm{d}x+\int_{\mathbb{R}}a\frac{\tau}{\mu}(\widetilde{\Pi}_{i})^{X_{i}}_{x}(\Pi-\widetilde{\Pi})\mathrm{d}x.$
We rewrite the function $Y_{i}$ as follows:
$Y_{i}:=\sum\limits_{j=1}\limits^{8}Y_{ij},$
where
$Y_{i1}:=\int_{\mathbb{R}}\frac{a}{\sigma_{i}}(\widetilde{u}_{i})^{X_{i}}_{x}(p(v)-p(\widetilde{v}))\mathrm{d}x,\quad Y_{i2}:=-\int_{\mathbb{R}}ap^{\prime}(\widetilde{v}_{i}^{X_{i}})(\widetilde{v}_{i})^{X_{i}}_{x}(v-\widetilde{v})\mathrm{d}x,$
$Y_{i3}:=\int_{\mathbb{R}}a(\widetilde{u}_{i})^{X_{i}}_{x}\left(u-\widetilde{u}-\frac{p(v)-p(\widetilde{v})}{\sigma_{i}}\right)\mathrm{d}x,\quad Y_{i4}:=-\int_{\mathbb{R}}a\left(p^{\prime}(\widetilde{v})-p^{\prime}(\widetilde{v}_{i}^{X_{i}})\right)(\widetilde{v}_{i})^{X_{i}}_{x}(v-\widetilde{v})\mathrm{d}x,$
$Y_{i5}:=\int_{\mathbb{R}}a\frac{\tau}{\mu}(\widetilde{\Pi}_{i})^{X_{i}}_{x}(\Pi-\widetilde{\Pi})\mathrm{d}x,\quad Y_{i6}:=-\int_{\mathbb{R}}(a_{i})_{x}^{X_{i}}\frac{\tau|\Pi-\widetilde{\Pi}|^{2}}{2\mu}\mathrm{d}x,$
$Y_{i7}:=-\frac{1}{2}\int_{\mathbb{R}}(a_{i})_{x}^{X_{i}}\left(u-\widetilde{u}-\frac{p(v)-p(\widetilde{v})}{\sigma_{i}}\right)\cdot\left(u-\widetilde{u}+\frac{p(v)-p(\widetilde{v})}{\sigma_{i}}\right)\mathrm{d}x,$
$Y_{i8}:=-\int_{\mathbb{R}}(a_{i})_{x}^{X_{i}}H(v|\widetilde{v})\mathrm{d}x-\frac{1}{2\sigma_{i}^{2}}\int_{\mathbb{R}}(a_{i})_{x}^{X_{i}}(p(v)-p(\widetilde{v}))^{2}\mathrm{d}x.$
Notice from (2.10) that
$\dot{X}_{i}=-\frac{M}{\delta_{i}}(Y_{i1}+Y_{i2}),$
which implies
$\dot{X}_{i}(t)Y_{i}(U)=\frac{\delta_{i}}{M}|\dot{X}_{i}|^{2}+\dot{X}_{i}\sum\limits_{j=3}\limits^{8}Y_{ij}.$ (4.7)
Next, we present the following lemma whose proof parallels the process outlined in Section 4.6 of reference [10]. Despite this methodological resemblance, we have observed that the approach in [10] employs cutoff functions to confine the perturbation, thereby obtaining dissipation estimates, which we have not employed in our proof. To preserve the integrity of our work and to illustrate the distinctions and relationships between our methodology and that of [10], we will provide a detailed exposition of this proof process in the ensuing lemma.
Lemma 4.5.
There exists a constant $C>0$ (independent of $\tau$) such that
$-\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{2M}|\dot{X}_{i}|^{2}+B_{1}+B_{2}-G_{2}-\frac{3}{2}D$
$\leq C\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}\left(-|(\widetilde{v}_{i})^{X_{i}}_{x}||p(v)-p(\widetilde{v})|^{2}+|(a_{i})^{X_{i}}_{x}||p(v)-p(\widetilde{v})|^{3}+|(a_{i})^{X_{i}}_{x}||\widetilde{v}-\widetilde{v}_{i}^{X_{i}}||p(v)-p(\widetilde{v})|^{2}\right)\mathrm{d}x$
where
$D=\int_{\mathbb{R}}\frac{a}{\gamma p(v)}|\partial_{x}(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x.$
Proof.
Let new variables $y_{1},y_{2}$ as follows:
$y_{1}:=\frac{p(\widetilde{v}_{1}(x-\sigma_{1}t))-p(v_{-})}{\delta_{1}},\qquad y_{2}:=\frac{p(v_{m})-p(\widetilde{v}_{2}(x-\sigma_{2}t))}{\delta_{2}}.$ (4.8)
For each $i$, using Lemma 2.1, we have
$\frac{\mathrm{d}y_{1}}{\mathrm{d}\xi_{1}}=\frac{1}{\delta_{1}}p^{\prime}(\widetilde{v}_{1})(\widetilde{v}_{1})_{\xi_{1}}>0,\quad\frac{\mathrm{d}y_{2}}{\mathrm{d}\xi_{2}}=-\frac{1}{\delta_{2}}p^{\prime}(\widetilde{v}_{2})(\widetilde{v}_{2})_{\xi_{2}}>0$
and
$\lim_{\xi_{i}\rightarrow-\infty}y_{i}=0,\quad\lim_{\xi_{i}\rightarrow+\infty}y_{i}=1.$
Then, we using new variables $y_{1},y_{2}$ to define new perturbation $w_{1},w_{2}$, respectively:
$w_{1}:=$ $\Big{(}p(v(t,x+X_{1}(t)))-p\left(\widetilde{v}_{1}(x-\sigma_{1}t)+\widetilde{v}_{2}(x-\sigma_{2}t-X_{2}(t)+X_{1}(t))-v_{m}\right)\Big{)}\circ y_{1}^{-1},$ (4.9)
$w_{2}:=$ $\Big{(}p(v(t,x+X_{2}(t)))-p\left(\widetilde{v}_{1}(x-\sigma_{1}t-X_{1}(t)+X_{2}(t))+\widetilde{v}_{2}(x-\sigma_{2}t)-v_{m}\right)\Big{)}\circ y_{2}^{-1}.$
In addition, the following estimates are hold:
$|\sigma_{1}-(-\sigma_{m})|<C\delta_{1},\quad|\sigma_{2}-\sigma_{m}|<C\delta_{2}$ (4.10)
and
$|\sigma_{m}^{2}-|p^{\prime}(\widetilde{v}_{i})||\leq C\delta_{i},\quad\Big{|}\frac{1}{\sigma_{m}^{2}}-\frac{p(\widetilde{v}_{i})^{-\frac{1}{\gamma}-1}}{\gamma}\Big{|}\leq C\delta_{i},$ $\Big{|}\frac{1}{\sigma_{m}^{2}}-\frac{p(\widetilde{v})^{-\frac{1}{\gamma}-1}}{\gamma}\Big{|}\leq C\delta_{0},$ (4.11)
where $\sigma_{m}=\sqrt{-p^{\prime}(v_{m})}$.
[…]
Using (2.2), we have
$Y_{11}=\frac{1}{\sigma_{1}^{2}}\int_{\mathbb{R}}a\left(p(\widetilde{v}_{1})^{X_{1}}_{x}-(\widetilde{\Pi}_{1})^{X_{1}}_{x}\right)(p(v)-p(\widetilde{v}))\mathrm{d}x.$
Changing the variable $x\rightarrow x+X_{1}(t)$ and applying the new variable for $y_{1},w_{1}$, we derive that
$\frac{1}{\sigma_{1}^{2}}\int_{\mathbb{R}}ap(\widetilde{v}_{1})^{X_{1}}_{x}(p(v)-p(\widetilde{v}))\mathrm{d}x=\frac{\delta_{1}}{\sigma_{1}^{2}}\int_{0}^{1}a(t,x+X_{1}(t))w_{1}\mathrm{d}y_{1}.$
Then, using (4.9), (4.11) and definition of $a(t,x)$ in (2.11), we have
$\Big{|}Y_{11}-\frac{\delta_{1}}{\sigma_{m}^{2}}\int_{0}^{1}w_{1}\mathrm{d}y_{1}\Big{|}\leq C\delta_{1}\delta_{0}\int_{0}^{1}|w_{1}|\mathrm{d}y_{1}+C\int_{\mathbb{R}}|(\widetilde{\Pi}_{1})^{X_{1}}_{x}||p(v)-p(\widetilde{v})|\mathrm{d}x.$
To estimates $Y_{12}$, we first notice that
$\Big{|}v-\widetilde{v}-\Big{(}-\frac{p(\widetilde{v})^{-\frac{1}{\gamma}-1}}{\gamma}(p(v)-p(\widetilde{v}))\Big{)}\Big{|}\leq C|p(v)-p(\widetilde{v})|^{2}.$
Then, using (3.5) and (4.11), we derive that
$\Big{|}v-\widetilde{v}-\Big{(}-\frac{1}{\sigma_{m}^{2}}(p(v)-p(\widetilde{v}))\Big{)}\Big{|}\leq C(\delta_{0}+\varepsilon_{1})|p(v)-p(\widetilde{v})|.$
Next, changing the variable $x\rightarrow x+X_{1}(t)$ and applying the new variable for $y_{1},w_{1}$, we derive that
$Y_{12}=-\delta_{1}\int_{0}^{1}a(t,x+X_{1}(t))(v-\widetilde{v})\mathrm{d}y_{1}.$
So, using (4.11) and the definition of $a(t,x)$ in (2.11), the above estimates implies
$\Big{|}Y_{12}-\frac{\delta_{1}}{\sigma_{m}^{2}}\int_{0}^{1}w_{1}\mathrm{d}y_{1}\Big{|}$ $\leq\delta_{1}\int_{0}^{1}a\Big{|}v-\widetilde{v}+\frac{w_{1}}{\sigma_{m}^{2}}\Big{|}\mathrm{d}y_{1}+\delta_{1}\int_{0}^{1}(a-1)\Big{|}\frac{w_{1}}{\sigma_{m}^{2}}\Big{|}\mathrm{d}y_{1}$
$\leq C\delta_{1}(\delta_{0}+\varepsilon_{1})\int_{0}^{1}|w_{1}|\mathrm{d}y_{1}.$
Combining the above estimates for $Y_{11}$ and $Y_{12}$, we get
$\Big{|}\dot{X}_{1}+\frac{2M}{\sigma_{m}^{2}}\int_{0}^{1}w_{1}\mathrm{d}y_{1}\Big{|}$ $\leq\frac{M}{\delta_{1}}\sum\limits_{j=1}\limits^{2}\Big{|}Y_{1j}-\frac{\delta_{1}}{\sigma_{m}^{2}}\int_{0}^{1}w_{1}\mathrm{d}y_{1}\Big{|}$
$\leq C(\delta_{0}+\varepsilon_{1})\int_{0}^{1}|w_{1}|\mathrm{d}y_{1}+\frac{C}{\delta_{1}}\int_{\mathbb{R}}|(\widetilde{\Pi}_{1})^{X_{1}}_{x}||p(v)-p(\widetilde{v})|\mathrm{d}x,$
which implies
$\left(\Big{|}\frac{2M}{\sigma_{m}^{2}}\int_{0}^{1}w_{1}\mathrm{d}y_{1}\Big{|}-|\dot{X}_{1}|\right)^{2}\leq C(\delta_{0}+\varepsilon_{1})^{2}\int_{0}^{1}|w_{1}|^{2}\mathrm{d}y_{1}+\frac{C}{\delta_{1}^{2}}\left(\int_{\mathbb{R}}|(\widetilde{\Pi}_{1})^{X_{1}}_{x}||p(v)-p(\widetilde{v})|\mathrm{d}x\right)^{2}.$
On the other hand, using Hölder inequality, Lemma 2.1 and using new variable $y_{1}$ and $w_{1}$, we have
$\frac{C}{\delta_{1}^{2}}\left(\int_{\mathbb{R}}|(\widetilde{\Pi}_{1})^{X_{1}}_{x}||p(v)-p(\widetilde{v})|\mathrm{d}x\right)^{2}$
$\leq\frac{C}{\delta_{1}^{2}}\int_{\mathbb{R}}|(\widetilde{\Pi}_{1})^{X_{1}}_{x}|\mathrm{d}x\int_{\mathbb{R}}|(\widetilde{\Pi}_{1})^{X_{1}}_{x}||(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x$
$\leq C\delta_{1}\int_{\mathbb{R}}|(\widetilde{v}_{1})^{X_{1}}_{x}||(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x$
$\leq C\delta_{1}^{2}\int_{0}^{1}|w_{1}|^{2}\mathrm{d}y_{1}.$
On the other hand, using the algebraic inequality, we derive that
$-\frac{\delta_{1}}{2M}|\dot{X}_{1}|^{2}\leq$ $-\frac{M\delta_{1}}{\sigma_{m}^{4}}\left(\int_{0}^{1}w_{1}\mathrm{d}y_{1}\right)^{2}+C\delta_{1}(\delta_{0}^{2}+\varepsilon_{1}^{2})\int_{0}^{1}|w_{1}|^{2}\mathrm{d}y_{1}.$
Next, we estimates the bad term $B_{1}$ and good term $G_{2}$. Recalling that
$B_{1}(U):=\sum\limits_{i=1}\limits^{2}\underbrace{\frac{1}{2\sigma_{i}}\int_{\mathbb{R}}(a_{i})_{x}^{X_{i}}(p(v)-p(\widetilde{v}))^{2}\mathrm{d}x}_{=:B_{1i}},\quad G_{2}(U):=\sum\limits_{i=1}\limits^{2}\underbrace{\sigma_{i}\int_{\mathbb{R}}(a_{i})^{X_{i}}_{x}H(v|\widetilde{v})}_{=:G_{2i}}\mathrm{d}x.$
For simplicity, we only estimates the case of $i=1$. Firstly, using Lemma 4.1, we obtain
$G_{21}\geq$ $\underbrace{\sigma_{1}\int_{\mathbb{R}}(a_{1})^{X_{1}}_{x}\frac{p(\widetilde{v}_{1}^{X_{1}})^{-\frac{1}{\gamma}-1}}{2\gamma}|p(v)-p(\widetilde{v})|^{2}\mathrm{d}x}_{=:\mathcal{G}_{1}}-\sigma_{1}\int_{\mathbb{R}}(a_{1})^{X_{1}}_{x}\frac{1+\gamma}{3\gamma}p(\widetilde{v})^{-\frac{1}{\gamma}-2}(p(v)-p(\widetilde{v}))^{3}\mathrm{d}x$
$+\frac{\sigma_{1}}{2\gamma}\int_{\mathbb{R}}(a_{1})^{X_{1}}_{x}\left(p(\widetilde{v})^{-\frac{1}{\gamma}-1}-p(\widetilde{v}_{1}^{X_{1}})^{-\frac{1}{\gamma}-1}\right)|p(v)-p(\widetilde{v})|^{2}\mathrm{d}x.$
Using (4.10) and (4.11), we have
$B_{11}\leq\frac{1}{2\sigma_{m}}\int_{\mathbb{R}}|(a_{1})^{X_{1}}_{x}||p(v)-p(\widetilde{v})|^{2}\mathrm{d}x+\frac{C\delta_{1}}{2\sigma_{m}\sigma_{1}}\int_{\mathbb{R}}|(a_{1})^{X_{1}}_{x}||p(v)-p(\widetilde{v})|^{2}\mathrm{d}x$
and
$\frac{p(\widetilde{v}_{1}^{X_{1}})^{-\frac{1}{\gamma}-1}}{2\gamma}\geq\frac{1}{2\sigma_{m}^{2}}-C\delta_{1},\quad-\frac{\sigma_{1}}{\sigma_{m}}\geq 1-C\delta_{1}.$
Hence, we have
$\mathcal{G}_{1}$ $\geq(\sigma_{m}-C\sigma_{m}\delta_{1})\left(\frac{1}{2\sigma_{m}^{2}}-C\delta_{1}\right)\int_{\mathbb{R}}|(a_{1})^{X_{1}}_{x}||p(v)-p(\widetilde{v})|^{2}\mathrm{d}x$
$\geq\frac{1}{2\sigma_{m}}(1-C\delta_{1})\int_{\mathbb{R}}|(a_{1})^{X_{1}}_{x}||p(v)-p(\widetilde{v})|^{2}\mathrm{d}x.$
Then, we derive that
$B_{11}-\mathcal{G}_{1}$ $\leq C\delta_{1}\int_{\mathbb{R}}|(a_{1})^{X_{1}}_{x}||p(v)-p(\widetilde{v})|^{2}\mathrm{d}x$
$\leq C\delta_{1}\lambda_{1}\int_{\mathbb{R}}\frac{|p((\widetilde{v}_{1})^{X_{1}}_{x})|}{\delta_{1}}|(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x.$
By change of variable $x\rightarrow x+X_{1}(t)$ and use the new variables $y_{1}$ and $w_{1}$, we get
$C\delta_{1}\lambda_{1}\int_{\mathbb{R}}\frac{|p((\widetilde{v}_{1})^{X_{1}}_{x})|}{\delta_{1}}|(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x=C\delta_{1}\lambda_{1}\int_{0}^{1}|w_{1}|^{2}\mathrm{d}y_{1}.$
Therefore, we have
$B_{11}-\mathcal{G}_{1}\leq C\delta_{1}\lambda_{1}\int_{0}^{1}|w_{1}|^{2}\mathrm{d}y_{1}.$
Now, we estimates the bad term $B_{2}$, which can be written as
$B_{2}(U):=\sum\limits_{i=1}\limits^{2}\underbrace{\sigma_{i}\int_{\mathbb{R}}a(\widetilde{v}_{i})^{X_{i}}_{x}p(v|\widetilde{v})\mathrm{d}x}_{=:B{2i}}.$
Similarly, we only estimates the case of $i=1$. Using (2.11), (4.8), (4.10), (4.11) and Lemma 4.1, we have
$B_{21}$ $=\sigma_{1}\delta_{1}\int_{0}^{1}a(t,x+X_{1}(t))\frac{1}{p^{\prime}(\widetilde{v}_{1})}p(v|\widetilde{v})^{-X_{1}}\mathrm{d}y_{1}$
$\leq|\sigma_{1}|\delta_{1}\int_{0}^{1}(1+C\delta_{0})\Big{|}\frac{1}{p^{\prime}(\widetilde{v}_{1})}\Big{|}\left(\frac{\gamma+1}{2\gamma\sigma_{m}p(v_{m})}\frac{\sigma_{m}p(v_{m})}{p(\widetilde{v})}+C\varepsilon_{1}\right)|w_{1}|^{2}\mathrm{d}y_{1}$
$\leq\delta_{1}(\sigma_{m}+C\delta_{0})(1+C\delta_{0})\left(\frac{1}{\sigma_{m}^{2}}+C\delta_{0}\right)\alpha_{m}\sigma_{m}(1+C(\delta_{0}+\varepsilon_{1}))\int_{0}^{1}|w_{1}|^{2}\mathrm{d}y_{1}$
$\leq\delta_{1}\alpha_{m}(1+C(\delta_{0}+\varepsilon_{1}))\int_{0}^{1}|w_{1}|^{2}\mathrm{d}y_{1},$
where $\alpha_{m}=\frac{\gamma+1}{2\gamma\sigma_{m}p(v_{m})}$.
On the other hand, define
$D(U):=\int_{\mathbb{R}}\frac{a}{\gamma p(v)}|\partial_{x}(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x.$
Using the methods in Lemma 4.5 of [25], we can get
$-D(U)\leq-\sum\limits_{i=1}\limits^{2}\delta_{i}\alpha_{m}(1-C(\delta_{0}+\varepsilon_{1}))\int_{0}^{1}y_{i}(1-y_{i})|\partial_{y_{i}}w_{i}|^{2}\mathrm{d}y_{i}.$
Combining the estimates of $\frac{\delta_{i}}{2M}|\dot{X}_{i}|^{2},B_{1},\mathcal{G}_{i},B_{2}$, we have
$B_{1}+B_{2}-\mathcal{G}-\frac{3}{2}D$
$\leq\sum\limits_{i=1}\limits^{2}\delta_{i}\alpha_{m}\left((1+C(\delta_{0}+\varepsilon_{1}))\int_{0}^{1}|w_{i}|^{2}\mathrm{d}y_{i}-\frac{3}{4}(1-C_{0}(\delta_{0}+\varepsilon_{1}))\int_{0}^{1}y_{i}(1-y_{i})|\partial_{y_{i}}w_{i}|^{2}\mathrm{d}y_{i}\right),$
where $\mathcal{G}:=\mathcal{G}_{1}+\mathcal{G}_{2}$ and
$\mathcal{G}_{i}:=\sigma_{i}\int_{\mathbb{R}}(a_{i})^{X_{i}}_{x}\frac{p(\widetilde{v}_{i}^{X_{i}})^{-\frac{1}{\gamma}-1}}{2\gamma}|p(v)-p(\widetilde{v})|^{2}\mathrm{d}x.$
In the following , we shall use the Poincaré-type inequality:(see [25])
$\int_{0}^{1}\Big{|}f-\int_{0}^{1}f\mathrm{d}y\Big{|}^{2}\mathrm{d}y\leq\int_{0}^{1}\frac{1}{2}y(1-y)|f^{\prime}|^{2}\mathrm{d}y,$
where $f:[0,1]\rightarrow\mathbb{R}$ with $\int_{0}^{1}y(1-y)|f^{\prime}|^{2}\mathrm{d}y<\infty$. Noting that
$\int_{0}^{1}|w-\bar{w}|^{2}\mathrm{d}y=\int_{0}^{1}w^{2}\mathrm{d}y-\bar{w}^{2},\quad\bar{w}:=\int_{0}^{1}w\mathrm{d}y,$
and using the smallness of $\delta_{0},\delta_{1},\delta_{2},\varepsilon_{1}$, we derive that
$B_{1}+B_{2}-\mathcal{G}-\frac{3}{4}D$
$\leq\sum\limits_{i=1}\limits^{2}\left(-\frac{\delta_{i}\alpha_{m}}{8}\int_{0}^{1}|w_{i}|^{2}\mathrm{d}y_{i}+\frac{5\delta_{i}\alpha_{m}}{4}\left(\int_{0}^{1}w_{i}\mathrm{d}y_{i}\right)^{2}\right).$
Finally, combining the estimates of $\frac{\delta_{i}}{2M}|\dot{X}_{i}|^{2}$ and choosing $M=\frac{5}{4}\sigma_{m}^{4}\alpha_{m}$, we have
$\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{2M}|\dot{X}_{i}|^{2}+B_{1}+B_{2}-G_{2}-\frac{3}{2}D$
$\leq\sum\limits_{i=1}\limits^{2}\left(-\frac{\delta_{i}\alpha_{m}}{16}\int_{0}^{1}|w_{i}|^{2}\mathrm{d}y_{i}-\frac{\sigma_{i}}{2\gamma}\int_{\mathbb{R}}(a_{i})^{X_{i}}_{x}\left(p(\widetilde{v})^{-\frac{1}{\gamma}-1}-p(\widetilde{v}_{i}^{X_{i}})^{-\frac{1}{\gamma}-1}|p(v)-p(\widetilde{v})|^{2}\mathrm{d}x\right)\right.$
$\qquad\left.+\sigma_{i}\int_{\mathbb{R}}(a_{i})^{X_{i}}_{x}\frac{1+\gamma}{3\gamma}p(\widetilde{v})^{-\frac{1}{\gamma}-2}(p(v)-p(\widetilde{v}))^{3}\mathrm{d}x\right)$
which implies
$\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{2M}|\dot{X}_{i}|^{2}+B_{1}+B_{2}-G_{2}-\frac{3}{2}D$
$\leq C\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}\left(-|(\widetilde{v}_{i})^{X_{i}}_{x}||(p(v)-p(\widetilde{v}))|^{2}+|(a_{i})^{X_{i}}_{x}||p(v)-p(\widetilde{v})|^{3}+|(a_{i})^{X_{i}}_{x}||\widetilde{v}-\widetilde{v}_{i}^{X_{i}}||p(v)-p(\widetilde{v})|^{2}\right)\mathrm{d}x.$
Thus, the proof of this lemma is finished. ∎
Now, we are ready to prove Lemma 4.2. Firstly, using Young’s inequality, we have
$\sum\limits_{i=1}\limits^{2}\left(\dot{X}_{i}\sum\limits_{j=3}\limits^{8}Y_{ij}\right)\leq\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{4M}|\dot{X}_{i}|^{2}+\sum\limits_{i=1}\limits^{2}\frac{C}{\delta_{i}}\sum\limits_{j=3}\limits^{8}|Y_{ij}|^{2}.$
Then, we derive that
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}}a\eta(U|\widetilde{U})\mathrm{d}x$
$\leq C(-\mathcal{G}^{s}+K_{1}+K_{2})-\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{4M}|\dot{X}_{i}|^{2}+\sum\limits_{i=1}\limits^{2}\frac{C}{\delta_{i}}\sum\limits_{j=3}\limits^{8}|Y_{ij}|^{2}+\sum\limits_{i=3}\limits^{7}B_{i}-G_{1}-G_{3}-G+\frac{3}{2}D,$
where
$\mathcal{G}^{s}:=\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x,\quad K_{1}:=\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}|(a_{i})^{X_{i}}_{x}||p(v)-p(\widetilde{v})|^{3}\mathrm{d}x,$
$K_{2}:=\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}|(a_{i})^{X_{i}}_{x}||\widetilde{v}-\widetilde{v}_{i}^{X_{i}}||p(v)-p(\widetilde{v})|^{2}\mathrm{d}x.$
For $K_{1}$, using the interpolation inequality, Hölder inequality, Young’s inequality, (2.13), Lemma 2.1 and $\lambda_{i}\leq C\sqrt{\delta_{i}}$, we have
$K_{1}$ $\leq\sum\limits_{i=1}\limits^{2}\frac{C\lambda_{i}}{\delta_{i}}\|p(v)-p(\widetilde{v})\|^{2}_{L^{\infty}}\sqrt{\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x}\sqrt{\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}|\mathrm{d}x}$
$\leq\sum\limits_{i=1}\limits^{2}\frac{C\lambda_{i}}{\sqrt{\delta_{i}}}\|\partial_{x}(p(v)-p(\widetilde{v}))\|_{L^{2}}\|p(v)-p(\widetilde{v})\|_{L^{2}}\sqrt{\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x}$
$\leq C\varepsilon_{1}(D+\mathcal{G}^{s}).$
Similarly, for $K_{2}$, by addtionally using Lemma 4.3, one obtain
$K_{2}$ $\leq\sum\limits_{i=1}\limits^{2}\frac{C\lambda_{i}}{\delta_{i}}\|p(v)-p(\widetilde{v})\|_{L^{4}}^{2}\left\||(\widetilde{v}_{i})^{X_{i}}_{x}||\widetilde{v}-\widetilde{v}_{i}^{X_{i}}|\right\|_{L^{2}}$
$\leq\sum\limits_{i=1}\limits^{2}\frac{C\lambda_{i}}{\delta_{i}}\|p(v)-p(\widetilde{v})\|_{L^{2}}^{3/2}\|\partial_{x}(p(v)-p(\widetilde{v}))\|_{L^{2}}^{1/2}\left\||(\widetilde{v}_{i})^{X_{i}}_{x}||\widetilde{v}-\widetilde{v}_{i}^{X_{i}}|\right\|_{L^{2}}$
$\leq C\varepsilon_{1}\sum\limits_{i=1}\limits^{2}\sqrt{\delta_{i}}\|\partial_{x}(p(v)-p(\widetilde{v}))\|_{L^{2}}^{1/2}\left\|\sqrt{|(\widetilde{v}_{i})^{X_{i}}_{x}|}|\widetilde{v}-\widetilde{v}_{i}^{X_{i}}|\right\|_{L^{2}}$
$\leq C\varepsilon_{1}\left(D+\sum\limits_{i=1}\limits^{2}\delta_{i}\left\|\sqrt{|(\widetilde{v}_{i})^{X_{i}}_{x}|}|\widetilde{v}-\widetilde{v}_{i}^{X_{i}}|\right\|_{L^{2}}^{2}\right).$
For $Y_{ij}$, using (2.13), Young’s inequality and Hölder inequality, we first get
$Y_{i3}$ $\leq C\frac{\delta_{i}}{\lambda_{i}}\int_{\mathbb{R}}|(a_{i})^{X_{i}}_{x}|\Big{|}u-\widetilde{u}-\frac{p(v)-p(\widetilde{v})}{\sigma_{i}}\Big{|}\mathrm{d}x$
$\leq C\frac{\delta_{i}}{\lambda_{i}}\left(\int_{\mathbb{R}}|(a_{i})^{X_{i}}_{x}|\mathrm{d}x\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}|(a_{i})^{X_{i}}_{x}|\Big{|}u-\widetilde{u}-\frac{p(v)-p(\widetilde{v})}{\sigma_{i}}\Big{|}^{2}\mathrm{d}x\right)^{\frac{1}{2}}$
$\leq C\frac{\delta_{i}}{\sqrt{\lambda_{i}}}\left(\int_{\mathbb{R}}|(a_{i})^{X_{i}}_{x}|\Big{|}u-\widetilde{u}-\frac{p(v)-p(\widetilde{v})}{\sigma_{i}}+\frac{\Pi-\widetilde{\Pi}}{\sigma_{i}}\Big{|}^{2}\mathrm{d}x+\int_{\mathbb{R}}|(a_{i})^{X_{i}}_{x}|\Big{|}\frac{\Pi-\widetilde{\Pi}}{\sigma_{i}}\Big{|}^{2}\mathrm{d}x\right)^{\frac{1}{2}}$
$\leq C\frac{\delta_{i}}{\sqrt{\lambda_{i}}}\sqrt{G_{1}+G}.$
Meanwhile, using Lemmas 2.1, 4.3 and Hölder inequality, we have
$Y_{i4}\leq C\int_{\mathbb{R}}|\widetilde{v}-\widetilde{v}_{i}^{X_{i}}||(\widetilde{v}_{i})^{X_{i}}_{x}||p(v)-p(\widetilde{v})|\mathrm{d}x$ $\leq C\delta_{0}\sqrt{\delta_{i}}\sqrt{\mathcal{G}^{s}}.$
Similarly, for $Y_{i5},Y_{i6},Y_{i7}$, we have
$Y_{i5}\leq C\tau\int_{\mathbb{R}}|(\widetilde{\Pi}_{i})^{X_{i}}_{x}||\Pi-\widetilde{\Pi}|\mathrm{d}x\leq C\tau\delta_{i}\sqrt{G},$
$Y_{i6}\leq\sqrt{\tau}\|\Pi-\widetilde{\Pi}\|_{L^{\infty}}\int_{\mathbb{R}}|(a_{i})_{x}^{X_{i}}|\frac{\sqrt{\tau}|\Pi-\widetilde{\Pi}|}{2\mu}\mathrm{d}x\leq C\varepsilon_{1}\lambda_{i}\sqrt{\tau\delta_{i}}\sqrt{G},$
and
$Y_{i7}\leq C\varepsilon_{1}\|(a_{i})_{x}^{X_{i}}\|_{L^{\infty}}^{\frac{1}{2}}\sqrt{G_{1}+G}\leq C\varepsilon_{1}\sqrt{\lambda_{i}\delta_{i}}\sqrt{G_{1}+G}.$
For $Y_{i8}$, by using (2.13), Lemmas 2.1, 4.3 and Hölder inequality, we get
$\frac{C}{\delta_{i}}|Y_{i8}|^{2}$ $\leq\frac{C\lambda_{i}^{2}}{\delta_{i}^{3}}\left(\int_{\mathbb{R}}|(\widetilde{v}_{i})_{x}^{X_{i}}||p(v)-p(\widetilde{v})|^{2}\mathrm{d}x\right)^{2}$
$\leq\frac{C\lambda_{i}^{2}}{\delta_{i}}\int_{\mathbb{R}}|p(v)-p(\widetilde{v})|^{2}\mathrm{d}x\int_{\mathbb{R}}|(\widetilde{v}_{i})_{x}^{X_{i}}||p(v)-p(\widetilde{v})|^{2}\mathrm{d}x$
$\leq C\varepsilon_{1}^{2}\mathcal{G}^{s}.$
Likewise, for $D$, we have
$D$ $\leq C\int_{\mathbb{R}}|\partial_{x}(v-\widetilde{v})|^{2}\mathrm{d}x+C\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}|(\widetilde{v}_{i})_{x}^{X_{i}}|^{2}|p(v)-p(\widetilde{v})|^{2}\mathrm{d}x$
$\leq C_{2}\underbrace{\int_{\mathbb{R}}|\partial_{x}(v-\widetilde{v})|^{2}\mathrm{d}x}_{:=\mathcal{D}}++C(\delta_{1}^{2}+\delta_{2}^{2})\mathcal{G}^{s}.$
For $B_{3}$, using $\|(a_{i})_{x}^{X_{i}}\|_{L^{\infty}}\leq C\lambda_{i}\delta_{i}$ and Lemma 2.1, we first have
$B_{3}\leq C(\lambda_{1}\delta_{1}+\lambda_{2}\delta_{2})G.$
For $B_{4}$, using (2.13), Lemma 2.1 and Young’s inequality, we have
$B_{4}\leq C\sum\limits_{i=1}\limits^{2}\left(\lambda_{i}\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}|^{\frac{1}{2}}|p(v)-p(\widetilde{v})||\Pi-\widetilde{\Pi}|\mathrm{d}x\right)\leq C(\lambda_{1}+\lambda_{2})\left(\mathcal{G}^{s}+G\right).$
Similarly, for $B_{5}$, using (3.1), $|\widetilde{\Pi}_{i}^{X_{i}}|\sim|(\widetilde{v}_{i})^{X_{i}}_{x}|$ and Young’s inequality, we have
$B_{5}$ $\leq C\int_{\mathbb{R}}(|(\widetilde{v}_{1})^{X_{1}}_{x}|+|(\widetilde{v}_{2})^{X_{2}}_{x}|)|\Pi-\widetilde{\Pi}||p(v)-p(\widetilde{v})|\mathrm{d}x$
$\leq C\delta_{i}\sum\limits_{i=1}\limits^{2}\left(\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||p(v)-p(\widetilde{v})|^{2}\mathrm{d}x+G\right)$
$\leq C(\delta_{1}+\delta_{2})\left(\mathcal{G}^{s}+G\right).$
For $B_{6},B_{7}$, we note that, for $i=1,2$
$F_{i}\leq C\left(|(\widetilde{v}_{1})^{X_{1}}_{x}||\widetilde{v}-(\widetilde{v}_{1})^{X_{1}}_{x}|+|(\widetilde{v}_{2})^{X_{2}}_{x}||\widetilde{v}-(\widetilde{v}_{2})^{X_{2}}_{x}|\right).$ (4.12)
So, using Lemma 4.3, Hölder inequality and Young’s inequality, we derive that
$|B_{6}|+|B_{7}|$ $\leq C\varepsilon_{1}\sum\limits_{i=1}\limits^{2}\left\||(\widetilde{v}_{i})^{X_{i}}_{x}||\widetilde{v}-(\widetilde{v}_{i})^{X_{i}}_{x}|\right\|_{L^{2}}+\frac{1}{64}G+C\sum\limits_{i=1}\limits^{2}\left\||(\widetilde{v}_{i})^{X_{i}}_{x}||\widetilde{v}-(\widetilde{v}_{i})^{X_{i}}_{x}|\right\|_{L^{2}}^{2}$
$\leq\frac{1}{64}G+C\varepsilon_{1}\delta_{1}\delta_{2}(\delta_{1}^{\frac{1}{2}}+\delta_{2}^{\frac{1}{2}})e^{-C\min\{\delta_{1},\delta_{2}\}t}+C\delta_{1}^{2}\delta_{2}^{2}(\delta_{1}+\delta_{2})e^{-C\min\{\delta_{1},\delta_{2}\}t}$
$\leq\frac{1}{64}G+C\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}.$
Combining all the above estimates, we derive that
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}}a\eta(U|\widetilde{U})\mathrm{d}x+\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{4M}|\dot{X}_{i}|^{2}+\frac{1}{2}G_{1}+G_{3}+\frac{1}{4}G+\frac{1}{2}\mathcal{G}^{s}$
$\leq\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}+C_{2}\mathcal{D}.$
Finally, integrating it over $[0,T]$ for any $t\leq T$, we derive that
$\int_{\mathbb{R}}a(t,x)\eta(U(t,x)|\widetilde{U}(t,x))\mathrm{d}x+\int_{0}^{t}\left(\sum\limits_{i=1}\limits^{2}\delta_{i}|\dot{X}_{i}|^{2}+G_{1}+G_{3}+G+\mathcal{G}^{s}\right)\mathrm{d}t$
$\leq C\int_{\mathbb{R}}a(0,x)\eta(U(0,x)|\widetilde{U}(0,x))\mathrm{d}x+C\delta_{0}+C_{2}\int_{0}^{t}\mathcal{D}\mathrm{d}t.$
Using Lemma 4.1, we have finished proof of Lemma 4.2
4.2. High-order energy estimates
In this section, we show the high-order energy estimates for the system (4.1). Let $\Phi=v-\widetilde{v},\Psi=u-\widetilde{u},Q=\Pi-\widetilde{\Pi}$, then we have the following lemma.
Lemma 4.6.
Under the hypotheses of Proposition 3.1, there exists $C>0$ (independent of $\tau$ and $T$) such that for all $t\in(0,T]$, we have
$\|\left(\partial_{x}\Phi,\partial_{x}\Psi\right)\|_{H^{1}}^{2}+\tau\|\partial_{x}Q\|_{H^{1}}^{2}+\int_{0}^{t}\|\partial_{x}Q\|_{H^{1}}^{2}\mathrm{d}t\leq C(\|\left(\partial_{x}\Phi_{0},\partial_{x}\Psi_{0}\right)\|_{H^{1}}^{2}+\tau\|\partial_{x}Q_{0}\|_{H^{1}}^{2})$ (4.13)
$\qquad+C(\delta_{1}+\delta_{2})\int_{0}^{t}G^{s}(U)\mathrm{d}t+C(\delta_{1}+\delta_{2}+\varepsilon_{1})\int_{0}^{t}\left(\|\left(\partial_{x}\Phi,\partial_{x}\Psi\right)\|_{H^{1}}^{2}+G(U)\right)\mathrm{d}t$
$\qquad+C\int_{0}^{t}\sum\limits_{i=1}^{2}\delta_{i}^{2}|\dot{X}_{i}|^{2}\mathrm{d}t+C\delta_{0},$
where $\Phi_{0}=v_{0}-\widetilde{v}_{0}(\xi),\Psi_{0}=u_{0}-\widetilde{u}_{0}(\xi),Q_{0}=\Pi_{0}-\widetilde{\Pi}_{0}(\xi)$, and $G^{s}(U),G(U)$ are defined in (3.7) and (4.4), respectively.
Proof.
Applying $\partial^{k}_{x}(k=1,2)$ to the system (4.1), we derive that
$\begin{cases}\partial_{t}\partial_{x}^{k}\Phi-\sum\limits_{i=1}\limits^{2}\dot{X}_{i}\partial_{x}^{k+1}\widetilde{v}_{i}^{X_{i}}-\partial_{x}^{k+1}\Psi=0,\\ \partial_{t}\partial_{x}^{k}\Psi-\sum\limits_{i=1}\limits^{2}\dot{X}_{i}\partial_{x}^{k+1}\widetilde{u}_{i}^{X_{i}}+p^{\prime}(v)\partial_{x}^{k+1}\Phi=\partial_{x}^{k+1}Q+F_{3}^{k},\\ \tau\partial_{t}\partial_{x}^{k}Q-\tau\sum\limits_{i=1}\limits^{2}\dot{X}_{i}\partial_{x}^{k+1}\widetilde{\Pi}_{i}^{X_{i}}+v\partial_{x}^{k}Q=\mu\partial_{x}^{k+1}\Psi+F_{4}^{k},\end{cases}$ (4.14)
where
$F_{3}^{k}=p^{\prime}(v)\partial^{k+1}_{x}(v-\widetilde{v})-\partial^{k+1}_{x}(p(v)-p(\widetilde{v}))-\partial^{k}_{x}F_{1}$
and
$F_{4}^{k}=v\partial^{k}_{x}(\Pi-\widetilde{\Pi})-\partial^{k}_{x}(v\Pi-\widetilde{v}\widetilde{\Pi})-\partial^{k}_{x}F_{2}.$
Multiplying the above equations by $-p^{\prime}(v)\partial_{x}^{k}\Phi,\partial_{x}^{k}\Psi,\frac{1}{\mu}\partial_{x}^{k}Q$, respectively, and integrating over $\mathbb{R}$, we get
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}}\left(-\frac{p^{\prime}(v)}{2}(\partial_{x}^{k}\Phi)^{2}+\frac{1}{2}(\partial_{x}^{k}\Psi)^{2}+\frac{\tau}{2\mu}(\partial_{x}^{k}Q)^{2}\right)\mathrm{d}x+\int_{\mathbb{R}}\frac{v}{\mu}(\partial_{x}^{k}Q)^{2}\mathrm{d}x=:\sum\limits_{i=1}\limits^{8}R^{k}_{i},$ (4.15)
where
$R^{k}_{1}=-\int_{\mathbb{R}}\frac{p^{\prime\prime}(v)}{2}v_{t}(\partial_{x}^{k}\Phi)^{2}\mathrm{d}x,\quad R^{k}_{2}=\int_{\mathbb{R}}p^{\prime\prime}(v)v_{x}\partial_{x}^{k}\Phi\partial_{x}^{k}\Psi\mathrm{d}x,$
$R^{k}_{3}=-\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}\dot{X}_{i}\partial_{x}^{k+1}\widetilde{v}^{X_{i}}_{i}p^{\prime}(v)\partial_{x}^{k}\Phi\mathrm{d}x,\quad R^{k}_{4}=\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}\dot{X}_{i}\partial_{x}^{k+1}\widetilde{u}^{X_{i}}_{i}\partial_{x}^{k}\Psi\mathrm{d}x,$
$R^{k}_{5}=\frac{\tau}{\mu}\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}\dot{X}_{i}\partial_{x}^{k+1}\widetilde{\Pi}^{X_{i}}_{i}Q\mathrm{d}x,\quad R^{k}_{6}=\int_{\mathbb{R}}F_{3}^{k}\partial_{x}^{k}\Psi\mathrm{d}x,\quad R^{k}_{7}=\frac{1}{\mu}\int_{\mathbb{R}}F_{4}^{k}\partial_{x}^{k}Q\mathrm{d}x.$
Firstly, using Lemmas 2.1 and Sobolev’s imbedding theorem, we have
$R^{k}_{1}=-\int_{\mathbb{R}}\frac{p^{\prime\prime}(v)}{2}(u-\widetilde{u})_{x}(\partial_{x}^{k}\Phi)^{2}\mathrm{d}x-\int_{\mathbb{R}}\frac{p^{\prime\prime}(v)}{2}\widetilde{u}_{x}(\partial_{x}^{k}\Phi)^{2}\mathrm{d}x\leq C(\varepsilon_{1}+\delta_{1}^{2}+\delta_{2}^{2})\int_{\mathbb{R}}(\partial_{x}^{k}\Phi)^{2}\mathrm{d}x.$
Similarly, for $R^{k}_{2}$, using Lemma 2.1, Sobolev’s imbedding theorem and Young’s inequality, we have
$R^{k}_{2}\leq C(\varepsilon_{1}+\delta_{1}^{2}+\delta_{2}^{2})\left(\int_{\mathbb{R}}(\partial_{x}^{k}\Phi)^{2}\mathrm{d}x+\int_{\mathbb{R}}(\partial_{x}^{k}\Psi)^{2}\mathrm{d}x\right).$
For $R^{k}_{3}$, using Lemma 2.1 and Young’s inequality, we have
$\int_{\mathbb{R}}\dot{X}_{i}\partial_{x}^{k+1}\widetilde{v}^{X_{i}}_{x}p^{\prime}(v)\partial_{x}^{k}\Phi\mathrm{d}x$ $\leq C\delta_{i}|\dot{X}_{i}|^{2}\int_{\mathbb{R}}|\partial_{x}^{k+1}\widetilde{v}^{X_{i}}_{i}|\mathrm{d}x+\frac{C}{\delta_{i}}\int_{\mathbb{R}}|\partial_{x}^{k+1}\widetilde{v}^{X_{i}}_{i}|(\partial_{x}^{k}\Phi)^{2}\mathrm{d}x$
$\leq C\delta_{i}^{2}|\dot{X}_{i}|^{2}+C\delta_{i}\int_{\mathbb{R}}(\partial_{x}^{k}\Phi)^{2}\mathrm{d}x.$
So, we get
$R^{k}_{3}\leq C\sum\limits_{i=1}\limits^{2}\left(\delta_{i}^{2}|\dot{X}_{i}|^{2}+\delta_{i}\int_{\mathbb{R}}(\partial_{x}^{k}\Phi)^{2}\mathrm{d}x\right).$
Similarly, for $R^{k}_{4},R^{k}_{5}$, we have
$R^{k}_{4}\leq C\sum\limits_{i=1}\limits^{2}\left(\delta_{i}^{2}|\dot{X}_{i}|^{2}+\delta_{i}\int_{\mathbb{R}}(\partial_{x}^{k}\Psi)^{2}\mathrm{d}x\right),\quad R^{k}_{5}\leq C\tau\sum\limits_{i=1}\limits^{2}\left(\delta_{i}^{2}|\dot{X}_{i}|^{2}+\delta_{i}\int_{\mathbb{R}}(\partial_{x}^{k}Q)^{2}\mathrm{d}x\right).$
Next, we estimate $R_{6}^{k}$. For $k=1$, since
$p^{\prime}(v)(v-\widetilde{v})_{xx}-(p(v)-p(\widetilde{v}))_{xx}$
$=-p^{\prime\prime}(v)v^{2}_{x}-p^{\prime}(v)\widetilde{v}_{xx}+p^{\prime\prime}(\widetilde{v})\widetilde{v}^{2}_{x}+p^{\prime}(\widetilde{v})\widetilde{v}_{xx}$
$=-p^{\prime\prime}(v)(v_{x}-\widetilde{v}_{x})^{2}-2p^{\prime\prime}(v)(v_{x}-\widetilde{v}_{x})\widetilde{v}_{x}-(p^{\prime\prime}(v)-p^{\prime\prime}(\widetilde{v}))\widetilde{v}^{2}_{x}-(p^{\prime}(v)-p^{\prime}(\widetilde{v}))\widetilde{v}_{xx},$
using Lemma 2.1, we have
$\int_{\mathbb{R}}(u-\widetilde{u})_{x}(p^{\prime}(v)(v-\widetilde{v})_{xx}-(p(v)-p(\widetilde{v}))_{xx})\mathrm{d}x$
$\leq C\int_{\mathbb{R}}|(u-\widetilde{u})_{x}||(v-\widetilde{v})_{x}|^{2}\mathrm{d}x+C\int_{\mathbb{R}}|(u-\widetilde{u})_{x}||(v-\widetilde{v})_{x}||\widetilde{v}_{x}|\mathrm{d}x+C\int_{\mathbb{R}}|(u-\widetilde{u})_{x}||v-\widetilde{v}||\widetilde{v}_{x}|\mathrm{d}x.$
We estimate each term of the right hand side of the above inequality. Using Young’s inequality and Sobolev’s imbedding theorem, we have
$\int_{\mathbb{R}}|(u-\widetilde{u})_{x}||(v-\widetilde{v})_{x}|^{2}\mathrm{d}x\leq C\varepsilon_{1}\int_{\mathbb{R}}|\partial_{x}\Phi|^{2}\mathrm{d}x$
and
$\int_{\mathbb{R}}|(u-\widetilde{u})_{x}||(v-\widetilde{v})_{x}||\widetilde{v}_{x}|\mathrm{d}x\leq C(\delta_{1}^{2}+\delta_{2}^{2})(\int_{\mathbb{R}}|\partial_{x}\Phi|^{2}\mathrm{d}x+\int_{\mathbb{R}}|\partial_{x}\Psi|^{2}\mathrm{d}x).$
In a similar way, using Lemma 2.1 and Young’s inequality, we get
$\int_{\mathbb{R}}|(u-\widetilde{u})_{x}||v-\widetilde{v}||(\widetilde{v}_{i})^{X_{i}}_{x}|\mathrm{d}x\leq C\delta_{i}\left(\int_{\mathbb{R}}|\partial_{x}\Psi|^{2}\mathrm{d}x+\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||v-\widetilde{v}|^{2}\mathrm{d}x\right).$
Hence, we derive that
$\int_{\mathbb{R}}(u-\widetilde{u})_{x}(p^{\prime}(v)(v-\widetilde{v})_{xx}-(p(v)-p(\widetilde{v}))_{xx})\mathrm{d}x$
$\leq C(\varepsilon_{1}+\delta_{1}+\delta_{2})\int_{\mathbb{R}}|\partial_{x}\Phi|^{2}\mathrm{d}x+C(\delta_{1}+\delta_{2})\int_{\mathbb{R}}|\partial_{x}\Psi|^{2}\mathrm{d}x+C(\delta_{1}+\delta_{2})G^{s}.$
On the other hand, using Lemma 2.1, we derive that for $k=1,2$,
$\partial_{x}^{k}F_{1}$ $=\partial_{x}^{k+1}\left(p(\widetilde{v})-p(\widetilde{v}_{1}^{X_{1}})-p(\widetilde{v}_{2}^{X_{2}})\right)$
$\leq C\left(|(\widetilde{v}_{1})^{X_{1}}_{x}||\widetilde{v}-(\widetilde{v}_{1})^{X_{1}}|+|(\widetilde{v}_{2})^{X_{2}}_{x}||\widetilde{v}-(\widetilde{v}_{2})^{X_{2}}|+|(\widetilde{v}_{1})^{X_{1}}_{x}||(\widetilde{v}_{2})^{X_{2}}_{x}|\right).$
So, using Lemmas 2.1, 4.3, Hölder inequality and Sobolev’s imbedding theorem, we have
-- $\int_{\mathbb{R}}(u-\widetilde{u})_{x}\partial_{x}F_{1}\mathrm{d}x$
$\leq C\varepsilon_{1}\left(\left\||(\widetilde{v}_{1})^{X_{1}}_{x}||\widetilde{v}-(\widetilde{v}_{1})^{X_{1}}|\right\|_{L^{2}}+\left\||(\widetilde{v}_{2})^{X_{2}}_{x}||\widetilde{v}-(\widetilde{v}_{2})^{X_{2}}|\right\|_{L^{2}}+\left\||(\widetilde{v}_{1})^{X_{1}}_{x}||(\widetilde{v}_{2})^{X_{2}}_{x}|\right\|_{L^{2}}\right)$
$\leq C\varepsilon_{1}\delta_{1}\delta_{2}(\delta_{1}^{\frac{1}{2}}+\delta_{2}^{\frac{1}{2}})e^{-C\min\{\delta_{1},\delta_{2}\}t}.$
Combining the above estimates, we derive that
$R_{6}^{1}$ $\leq C(\varepsilon_{1}+\delta_{1}+\delta_{2})\int_{\mathbb{R}}|\partial_{x}\Phi|^{2}\mathrm{d}x$
$+C(\delta_{1}+\delta_{2})\int_{\mathbb{R}}|\partial_{x}\Psi|^{2}\mathrm{d}x+C(\delta_{1}+\delta_{2})G^{s}+C\varepsilon_{1}\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}.$
Similarly, for $k=2$, since
$p^{\prime}(v)(v-\widetilde{v})_{xxx}-(p(v)-p(\widetilde{v}))_{xxx}$
$=-p^{\prime\prime\prime}(v)(v_{x}-\widetilde{v}_{x})^{3}-(p^{\prime\prime\prime}(v)-p^{\prime\prime\prime}(\widetilde{v}))\widetilde{v}_{x}^{3}-3p^{\prime\prime\prime}(v)\widetilde{v}_{x}^{2}(v_{\xi}-\widetilde{v}_{x})-3p^{\prime\prime\prime}(v)\widetilde{v}_{x}(v_{x}-\widetilde{v}_{x})^{2}$
$\quad-3p^{\prime\prime}(v)(v_{x}-\widetilde{v}_{x})(v_{xx}-\widetilde{v}_{xx})-3p^{\prime\prime}(v)\widetilde{v}_{x}(v_{xx}-\widetilde{v}_{xx})-3p^{\prime\prime}(v)\widetilde{v}_{xx}(v_{x}-\widetilde{v}_{x})$
$\quad-3(p^{\prime\prime}(v)-p^{\prime\prime}(\widetilde{v}))\widetilde{v}_{x}\widetilde{v}_{xx}-(p^{\prime}(v)-p^{\prime}(\widetilde{v}))\widetilde{v}_{xxx}.$
Using Lemma 2.1, Young’s inequality and Sobolev’s imbedding theorem, we have
$\int_{\mathbb{R}}(u-\widetilde{u})_{xx}(p^{\prime}(v)(v-\widetilde{v})_{xxx}-(p(v)-p(\widetilde{v}))_{xxx})\mathrm{d}x$
$\leq C\int_{\mathbb{R}}|(u-\widetilde{u})_{xx}||v-\widetilde{v}||\widetilde{v}_{x}|\mathrm{d}x+C\int_{\mathbb{R}}|(u-\widetilde{u})_{xx}||(v-\widetilde{v})_{x}|^{3}\mathrm{d}x$
$\quad+C\int_{\mathbb{R}}|(u-\widetilde{u})_{xx}||(v-\widetilde{v})_{x}||\widetilde{v}_{x}|\mathrm{d}x+C\int_{\mathbb{R}}|(u-\widetilde{u})_{xx}||(v-\widetilde{v})_{xx}||\widetilde{v}_{x}|\mathrm{d}x$
$\leq C(\varepsilon_{1}^{2}+\delta_{1}+\delta_{2})\left(\int_{\mathbb{R}}|\partial_{x}\Phi|^{2}\mathrm{d}x+\int_{\mathbb{R}}|\partial_{xx}\Psi|^{2}\mathrm{d}x\right)+C(\delta_{1}+\delta_{2})\int_{\mathbb{R}}|\partial_{xx}\Phi|^{2}\mathrm{d}x+C(\delta_{1}+\delta_{2})G^{s}$
and
-- $\int_{\mathbb{R}}(u-\widetilde{u})_{xx}\partial_{xx}F_{1}\mathrm{d}x$
$\leq C\varepsilon_{1}\left(\left\||(\widetilde{v}_{1})^{X_{1}}_{x}||\widetilde{v}-(\widetilde{v}_{1})^{X_{1}}_{x}|\right\|_{L^{2}}+\left\||(\widetilde{v}_{2})^{X_{2}}_{x}||\widetilde{v}-(\widetilde{v}_{2})^{X_{2}}_{x}|\right\|_{L^{2}}+\left\||(\widetilde{v}_{1})^{X_{1}}_{x}||(\widetilde{v}_{2})^{X_{2}}_{x}|\right\|_{L^{2}}\right)$
$\leq C\varepsilon_{1}\delta_{1}\delta_{2}(\delta_{1}^{\frac{1}{2}}+\delta_{2}^{\frac{1}{2}})e^{-C\min\{\delta_{1},\delta_{2}\}t}.$
So, we have
$R_{6}^{2}\leq C(\varepsilon_{1}^{2}+\delta_{1}+\delta_{2})\left(\int_{\mathbb{R}}|\partial_{x}\Phi|^{2}\mathrm{d}x+\int_{\mathbb{R}}|\partial_{xx}\Psi|^{2}\mathrm{d}x\right)+C(\delta_{1}+\delta_{2})\int_{\mathbb{R}}|\partial_{xx}\Phi|^{2}\mathrm{d}x$
$\qquad+C(\delta_{1}+\delta_{2})G^{s}+C\varepsilon_{1}\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}.$
Similarly, for $R_{7}^{1}$, since
$v(\Pi-\widetilde{\Pi})_{x}-(v\Pi-\widetilde{v}\widetilde{\Pi})_{x}=-(v-\widetilde{v})\widetilde{\Pi}_{x}-(v-\widetilde{v})_{x}(\Pi-\widetilde{\Pi})-(v-\widetilde{v})_{x}\widetilde{\Pi}-\widetilde{v}_{x}(\Pi-\widetilde{\Pi}),$
using Lemma 2.1 and $|\widetilde{\Pi}_{i}^{X_{i}}|\leq C|(\widetilde{v}_{i})^{X_{i}}_{x}|$, we have
$\int_{\mathbb{R}}\frac{(\Pi-\widetilde{\Pi})_{x}}{\mu}\left(v(\Pi-\widetilde{\Pi})_{x}-(v\Pi-\widetilde{v}\widetilde{\Pi})_{x}\right)\mathrm{d}x$
$\leq C\int_{\mathbb{R}}\frac{|(\Pi-\widetilde{\Pi})_{x}|}{\mu}|v-\widetilde{v}||\widetilde{v}_{x}|\mathrm{d}x+C\int_{\mathbb{R}}\frac{|(\Pi-\widetilde{\Pi})_{x}|}{\mu}|(v-\widetilde{v})_{x}||\Pi-\widetilde{\Pi}|\mathrm{d}x$
$\quad+C\int_{\mathbb{R}}\frac{|(\Pi-\widetilde{\Pi})_{x}|}{\mu}|(v-\widetilde{v})_{x}||\widetilde{\Pi}|\mathrm{d}x+C\int_{\mathbb{R}}\frac{|(\Pi-\widetilde{\Pi})_{x}|}{\mu}|\Pi-\widetilde{\Pi}||\widetilde{v}_{x}|\mathrm{d}x$
$\leq C(\varepsilon_{1}+\delta_{1}+\delta_{2})\left(\int_{\mathbb{R}}\frac{v}{\mu}(\partial_{x}Q)^{2}\mathrm{d}x+G\right)+(\delta_{1}+\delta_{2})\int_{\mathbb{R}}(\partial_{x}\Phi)^{2}\mathrm{d}x+(\delta_{1}+\delta_{2})G^{s}.$
On the other hand, for $k=1,2$, using Lemma 2.1, we get
$\partial_{x}^{k}F_{2}$ $=\partial_{x}^{k}\left(\left(\widetilde{v}^{X_{2}}_{2}-v_{m}\right)\widetilde{\Pi}_{1}^{X_{1}}+\left(\widetilde{v}^{X_{1}}_{1}-v_{m}\right)\widetilde{\Pi}_{2}^{X_{2}}\right)$
$\leq C\left(|(\widetilde{v}_{1})^{X_{1}}_{x}||\widetilde{v}-(\widetilde{v}_{1})^{X_{1}}_{x}|+|(\widetilde{v}_{2})^{X_{2}}_{x}||\widetilde{v}-(\widetilde{v}_{2})^{X_{2}}_{x}|+|(\widetilde{v}_{1})^{X_{1}}_{x}||(\widetilde{v}_{2})^{X_{2}}_{x}|\right).$
Then, using Lemma 4.3 and Young’s inequality, we have
$\int_{\mathbb{R}}\frac{(\Pi-\widetilde{\Pi})_{x}}{\mu}\left(\left(\widetilde{v}^{X_{2}}_{2}-v_{m}\right)\widetilde{\Pi}_{1}^{X_{1}}+\left(\widetilde{v}^{X_{1}}_{1}-v_{m}\right)\widetilde{\Pi}_{2}^{X_{2}}\right)_{x}\mathrm{d}x$
$\leq C\|\partial_{x}F_{2}\|_{L^{2}}+\frac{1}{16}\int_{\mathbb{R}}\frac{v}{\mu}(\partial_{x}Q)^{2}\mathrm{d}x$
$\leq C\delta_{1}^{2}\delta_{2}^{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}+\frac{1}{16}\int_{\mathbb{R}}\frac{v}{\mu}(\partial_{x}Q)^{2}\mathrm{d}x.$
Combining the above estimates, we derive that
$R_{7}^{1}\leq\frac{1}{8}\int_{\mathbb{R}}\frac{v}{\mu}(\partial_{x}Q)^{2}\mathrm{d}x+(\varepsilon_{1}+\delta_{1}+\delta_{2})G+(\delta_{1}+\delta_{2})\left(\int_{\mathbb{R}}(\partial_{x}\Phi)^{2}\mathrm{d}x+G^{s}\right)+C\delta_{1}^{2}\delta_{2}^{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}.$
Next, for $R_{7}^{2}$, first we have
$v(\Pi-\widetilde{\Pi})_{xx}-(v\Pi-\widetilde{v}\widetilde{\Pi})_{xx}=$ $-(v-\widetilde{v})_{xx}(\Pi-\widetilde{\Pi})-(v-\widetilde{v})_{xx}\widetilde{\Pi}-\widetilde{v}_{xx}(\Pi-\widetilde{\Pi})-(v-\widetilde{v})\widetilde{\Pi}_{xx}$
$-2(v-\widetilde{v})_{x}(\Pi-\widetilde{\Pi})_{x}-2(v-\widetilde{v})_{x}\widetilde{\Pi}_{x}-2\widetilde{v}_{x}(\Pi-\widetilde{\Pi})_{x}.$
Specially, using Hölder inequality, Sobolev’s imbedding theorem and Young’s inequality, we have
$\int_{\mathbb{R}}(v-\widetilde{v})_{xx}(\Pi-\widetilde{\Pi})(\Pi-\widetilde{\Pi})_{xx}\mathrm{d}x$
$\leq C\|(v-\widetilde{v})_{\xi\xi}\|_{L^{2}}\left(\int_{\mathbb{R}}\left(\frac{v}{\mu}\right)^{2}(\Pi-\widetilde{\Pi})^{2}(\Pi-\widetilde{\Pi})_{xx}^{2}\mathrm{d}x\right)^{\frac{1}{2}}$
$\leq C\varepsilon_{1}|\sqrt{\frac{v}{\mu}}(\Pi-\widetilde{\Pi})|_{L^{\infty}}\|\sqrt{\frac{v}{\mu}}(\Pi-\widetilde{\Pi})_{xx}\|_{L^{2}}$
$\leq C\varepsilon_{1}\left(\|\sqrt{\frac{v}{\mu}}(\Pi-\widetilde{\Pi})\|_{H^{1}}^{2}+\|\sqrt{\frac{v}{\mu}}(\Pi-\widetilde{\Pi})_{xx}\|_{L^{2}}^{2}\right).$
So, we have
$\int_{\mathbb{R}}\frac{(\Pi-\widetilde{\Pi})_{xx}}{\mu}\left(v(\Pi-\widetilde{\Pi})_{xx}-(v\Pi-\widetilde{v}\widetilde{\Pi})_{xx}\right)\mathrm{d}x$
$\leq C(\varepsilon_{1}+\delta_{1}+\delta_{2})\left\|\sqrt{\frac{v}{\mu}}Q\right\|_{H^{2}}+(\delta_{1}+\delta_{2})\left\|\partial_{x}\Phi\right\|_{H^{1}}+(\delta_{1}+\delta_{2})G^{s}.$
On the other hand, using the definition of $F_{2}$ and Young’s inequality, we have
$\int_{\mathbb{R}}\frac{(\Pi-\widetilde{\Pi})_{xx}}{\mu}\left(\left(\widetilde{v}^{X_{2}}_{2}-v_{m}\right)\widetilde{S}_{1}^{X_{1}}+\left(\widetilde{v}^{X_{1}}_{1}-v_{m}\right)\widetilde{\Pi}_{2}^{X_{2}}\right)_{xx}\mathrm{d}x$
$\leq C\|\partial_{xx}F_{2}\|_{L^{2}}+\frac{1}{16}\int_{\mathbb{R}}\frac{v}{\mu}(\partial_{xx}Q)^{2}\mathrm{d}x$
$\leq C\delta_{1}^{2}\delta_{2}^{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}+\frac{1}{16}\int_{\mathbb{R}}\frac{v}{\mu}(\partial_{xx}Q)^{2}\mathrm{d}x.$
Therefore, we get
$R_{7}^{2}$ $\leq\frac{1}{8}\int_{\mathbb{R}}\frac{v}{\mu}(\partial_{xx}Q)^{2}\mathrm{d}x$
$+C(\varepsilon_{1}+\delta_{1}+\delta_{2})\left\|\sqrt{\frac{v}{\mu}}Q\right\|_{H^{1}}+C(\delta_{1}+\delta_{2})\left(G^{s}+\left\|\partial_{x}\Phi\right\|_{H^{1}}\right)+C\delta_{1}^{2}\delta_{2}^{2}e^{-C\min\{\delta_{1},\delta_{2}\}t}.$
Finally, integrating the equality (4.15) over $[0,t]$, using the above estimates, we complete the proof of this lemma. ∎
4.3. Dissipative estimates
In the following lemmas, we give the dissipative estimates of given solutions to system (4.1).
Lemma 4.7.
Under the hypotheses of Proposition 3.1, there exist $C,C_{1},\nu>0$ (independent of $\tau$ and $T$) such that for all $t\in(0,T]$, we have
$\int_{0}^{t}\|\partial_{x}\Phi\|_{H^{1}}^{2}\mathrm{d}t\leq\nu\|\Psi\|_{H^{1}}^{2}+C(\nu)\|\partial_{x}\Phi\|_{H^{1}}^{2}+C\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}|\dot{X}_{i}(t)|^{2}\mathrm{d}t+C\int_{0}^{t}\|\partial_{x}Q\|_{H^{1}}^{2}\mathrm{d}t$ (4.16)
$+C\left(\|\Psi_{0}\|_{H^{1}}^{2}+\|\partial_{x}\Phi_{0}\|_{H^{1}}^{2}\right)+C_{1}\int_{0}^{t}\|\partial_{x}\Psi\|_{H^{1}}^{2}\mathrm{d}t+C(\delta_{1}+\delta_{2})\int_{0}^{t}G^{s}(U)\mathrm{d}t+C\delta_{0},$
where $G^{s}(U)$ is defined in (3.7).
Proof.
Multiplying the equation (4.14)2italic-(4.14subscriptitalic-)$_{2}$ by $\partial_{x}^{k+1}\Phi$ for $k=0,1,$ and integrating over $(0,t)\times\mathbb{R}$, we get
$\int_{0}^{t}\int_{\mathbb{R}}-p^{\prime}(v)\left(\partial_{x}^{k+1}\Phi\right)^{2}\mathrm{d}x\mathrm{d}t=:\sum\limits_{i=0}\limits^{4}M_{i}^{k},$
where
$M_{1}^{k}=\int_{0}^{t}\int_{\mathbb{R}}\partial_{t}\partial_{x}^{k}\Psi\partial_{x}^{k+1}\Phi\mathrm{d}x\mathrm{d}t,\quad M_{2}^{k}=-\sum\limits_{i=1}\limits^{2}\int_{0}^{t}\int_{\mathbb{R}}\dot{X_{i}}\partial_{x}^{k+1}\widetilde{u}_{i}^{X_{i}}\partial_{x}^{k+1}\Phi\mathrm{d}x\mathrm{d}t,$
$M_{3}^{k}=-\int_{0}^{t}\int_{\mathbb{R}}\partial_{x}^{k+1}Q\partial_{x}^{k+1}\Phi\mathrm{d}x\mathrm{d}t,\quad M_{4}^{k}=-\int_{0}^{t}\int_{\mathbb{R}}F_{3}^{k}\partial_{x}^{k+1}\Phi\mathrm{d}x\mathrm{d}t.$
[…]
$M_{1}^{k}=\int_{\mathbb{R}}\left(\partial_{x}^{k}\Psi(t)\partial_{x}^{k+1}\Phi(t)-\partial_{x}^{k}\Psi_{0}\partial_{x}^{k+1}\Phi_{0}\right)\mathrm{d}x-\int_{0}^{t}\int_{\mathbb{R}}\partial_{x}^{k}\Psi\left(\sum\limits_{i=1}\limits^{2}\dot{X}_{i}\partial_{x}^{k+2}\widetilde{v}_{i}^{X_{i}}+\partial_{x}^{k+2}\Psi\right)\mathrm{d}x\mathrm{d}t.$
Since
$-\int_{0}^{t}\int_{\mathbb{R}}\partial_{x}^{k}\Psi\left(\sum\limits_{i=1}\limits^{2}\dot{X}_{i}\partial_{x}^{k+2}\widetilde{v}_{i}^{X_{i}}\right)\mathrm{d}x\mathrm{d}t$ $=\int_{0}^{t}\int_{\mathbb{R}}\partial_{x}^{k+1}\Psi\left(\sum\limits_{i=1}\limits^{2}\dot{X}_{i}\partial_{x}^{k+1}\widetilde{v}_{i}^{X_{i}}\right)\mathrm{d}x\mathrm{d}t$
$\leq C\sum\limits_{i=1}\limits^{2}\left(\delta_{i}^{2}\int_{0}^{t}|\dot{X}_{i}|^{2}\mathrm{d}t+\delta_{i}\int_{0}^{t}\int_{\mathbb{R}}(\partial_{x}^{k+1}\Psi)^{2}\mathrm{d}x\mathrm{d}t\right)$
and
$-\int_{0}^{t}\int_{\mathbb{R}}\partial_{x}^{k}\Psi\partial_{x}^{k+2}\Psi\mathrm{d}x\mathrm{d}t=\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Psi\right)^{2}\mathrm{d}x\mathrm{d}t,$
we obtain
$M_{1}^{k}\leq$ $\nu\int_{\mathbb{R}}\left(\partial_{x}^{k}\Psi(t)\right)^{2}\mathrm{d}x+C(\nu)\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Phi(t)\right)^{2}\mathrm{d}x+C\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Psi\right)^{2}\mathrm{d}x\mathrm{d}t$
$+C\left(\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}\int_{0}^{t}|\dot{X}_{i}|^{2}\mathrm{d}t+\int_{\mathbb{R}}\left(\partial_{x}^{k}\Psi_{0}\right)^{2}\mathrm{d}x+\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Phi_{0}\right)^{2}\mathrm{d}x\right).$
Secondly, for $M_{2}^{k}$ and $M_{3}^{k}$, using Young’s inequality and Lemma 2.1, we have
$M_{2}^{k}\leq C\sum\limits_{i=1}\limits^{2}\left(\delta_{i}^{2}\int_{0}^{t}|\dot{X}_{i}|^{2}\mathrm{d}t+\delta_{i}\int_{0}^{t}\int_{\mathbb{R}}(\partial_{x}^{k+1}\Phi)^{2}\mathrm{d}x\mathrm{d}t\right)$
and
$M_{3}^{k}\leq\frac{1}{2}\int_{0}^{t}\int_{\mathbb{R}}-p^{\prime}(v)\left(\partial_{x}^{k+1}\Phi\right)^{2}\mathrm{d}x\mathrm{d}t+C\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}Q\right)^{2}\mathrm{d}x\mathrm{d}t.$
Next, we estimate $M_{4}^{k}$. For $k=0$, we have
$F_{3}^{0}=-(p^{\prime}(v)-p^{\prime}(\widetilde{v}))\widetilde{v}_{x}-\left(p(\widetilde{v})_{x}-p(\widetilde{v}_{1})^{X_{1}}_{x}-p(\widetilde{v}_{2})^{X_{2}}_{x}\right),$
which gives
$M_{4}^{0}\leq C(\delta_{1}+\delta_{2})\left(\int_{0}^{t}\int_{\mathbb{R}}|\partial_{x}\Phi|^{2}\mathrm{d}x\mathrm{d}t+\int_{0}^{t}G^{s}\mathrm{d}t\right)+C\varepsilon_{1}\max\{\delta_{1},\delta_{2}\}.$
For $k=1$, the term $F_{3}^{1}$ is the same as in Lemma 4.6. Therefore, using similar estimates as in Lemma 4.6, we have
$M_{4}^{1}$ $\leq C(\varepsilon_{1}+\delta_{1}+\delta_{2})\int_{0}^{t}\|\partial_{x}\Phi\|_{H^{1}}^{2}\mathrm{d}t+C(\delta_{1}+\delta_{2})\int_{0}^{t}G^{s}\mathrm{d}t+C\varepsilon_{1}\max\{\delta_{1},\delta_{2}\}.$
Therefore, combining the above estimates, we get the desired results. ∎
Lemma 4.8.
Under the hypotheses of Proposition 3.1, there exist $C,\nu_{1}>0$ (independent of $\tau$ and $T$) such that for all $t\in(0,T]$, we have
$\int_{0}^{t}\|\partial_{x}\Psi\|_{H^{1}}^{2}\mathrm{d}t\leq\nu_{1}\tau\|Q\|_{H^{1}}^{2}+C(\nu_{1})\|\partial_{x}\Psi\|_{H^{1}}^{2}+C\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}|\dot{X}_{i}(t)|^{2}\mathrm{d}t+\frac{1}{2C_{1}}\int_{0}^{t}\|\partial_{x}\Phi\|_{H^{1}}^{2}\mathrm{d}t$ (4.17)
$+C\left(\tau\|Q_{0}\|_{H^{1}}^{2}+\|\partial_{x}\Psi_{0}\|_{H^{1}}^{2}\right)+C\int_{0}^{t}\|\partial_{x}Q\|_{H^{1}}^{2}\mathrm{d}t+C(\varepsilon_{1}+\delta_{1}+\delta_{2})\int_{0}^{t}(G^{s}(U)+G(U))\mathrm{d}t+C\delta_{0},$
where $C_{1}$ is the constant in Lemma 4.7, and $G^{s}(U),G(U)$ are defined in (3.7) and (4.4), respectively.
Proof.
Multiplying the equation (4.14)3italic-(4.14subscriptitalic-)$_{3}$ by $\partial_{x}^{k+1}\Psi$ for $k=0,1$, and integrating over $(0,t)\times\mathbb{R}$, we get
$\int_{0}^{t}\int_{\mathbb{R}}\mu\left(\partial_{x}^{k+1}\Psi\right)^{2}\mathrm{d}x\mathrm{d}t=:\sum\limits_{i=0}\limits^{4}N_{i}^{k},$
where
$N_{1}^{k}=\int_{0}^{t}\int_{\mathbb{R}}\tau\partial_{t}\partial_{x}^{k}Q\partial_{x}^{k+1}\Psi\mathrm{d}x\mathrm{d}t,\quad N_{2}^{k}=-\tau\sum\limits_{i=1}\limits^{2}\dot{X}_{i}\partial_{x}^{k+1}\widetilde{\Pi}_{i}^{X_{i}}\partial_{x}^{k+1}\Psi\mathrm{d}x\mathrm{d}t,$
$N_{3}^{k}=\int_{0}^{t}\int_{\mathbb{R}}v\partial_{x}^{k}Q\partial_{x}^{k+1}\Psi\mathrm{d}x\mathrm{d}t,\quad N_{4}^{k}=-\int_{0}^{t}\int_{\mathbb{R}}F_{4}^{k}\partial_{x}^{k+1}\Psi\mathrm{d}x\mathrm{d}t.$
Firstly, doing integration by part and using equation (4.14)2italic-(4.14subscriptitalic-)$_{2}$, we get
$N_{1}^{k}=$ $\tau\int_{\mathbb{R}}\partial_{x}^{k}Q(t)\partial_{x}^{k+1}\Psi(t)\mathrm{d}x-\tau\int_{\mathbb{R}}\partial_{x}^{k}Q_{0}\partial_{x}^{k+1}\Psi_{0}\mathrm{d}x$
$-\tau\int_{0}^{t}\int_{\mathbb{R}}\partial_{x}^{k}Q\left(\sum\limits_{i=1}\limits^{2}\dot{X}_{i}\partial_{x}^{k+2}\widetilde{u}_{i}^{X_{i}}-\partial_{x}^{k+2}(p(v)-p(\widetilde{v}))+\partial_{x}^{k+2}Q+\partial_{x}^{k+2}F_{1}\right)\mathrm{d}x\mathrm{d}t.$
Integrating by part and using lemma 2.1 and Young’s inequality, we get
$-\tau\int_{0}^{t}\int_{\mathbb{R}}\partial_{x}^{k}Q\sum\limits_{i=1}\limits^{2}\dot{X}_{i}\partial_{x}^{k+2}\widetilde{u}_{i}^{X_{i}}\mathrm{d}x\mathrm{d}t$ $=\tau\int_{0}^{t}\int_{\mathbb{R}}\partial_{x}^{k+1}Q\sum\limits_{i=1}\limits^{2}\dot{X}_{i}\partial_{x}^{k+1}\widetilde{u}_{i}^{X_{i}}\mathrm{d}x\mathrm{d}t$
$\leq C\tau\sum\limits_{i=1}\limits^{2}\left(\delta_{i}^{2}\int_{0}^{t}|\dot{X}_{i}|^{2}\mathrm{d}t+\delta_{i}\int_{0}^{t}\int_{\mathbb{R}}(\partial_{x}^{k+1}Q)^{2}\mathrm{d}x\mathrm{d}t\right),$
$-\tau\int_{0}^{t}\int_{\mathbb{R}}\partial_{x}^{k}Q\partial_{x}^{k+2}Q\mathrm{d}x\mathrm{d}t=\tau\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}Q\right)^{2}\mathrm{d}x\mathrm{d}t,$
and
$\tau\int_{0}^{t}\int_{\mathbb{R}}\partial_{x}^{k}Q\partial_{x}^{k+2}(p(v)-p(\widetilde{v}))\mathrm{d}x\mathrm{d}t$
$=-\tau\int_{0}^{t}\int_{\mathbb{R}}\partial_{x}^{k+1}Q\partial_{x}^{k+1}(p(v)-p(\widetilde{v}))\mathrm{d}x\mathrm{d}t$
$\leq\frac{\mu}{4C_{1}}\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Phi\right)^{2}\mathrm{d}x\mathrm{d}t+C\tau^{2}\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}Q\right)^{2}\mathrm{d}x\mathrm{d}t+C(\delta_{1}+\delta_{2})\int_{0}^{t}G^{s}\mathrm{d}t+C\max\{\delta_{1},\delta_{2}\}.$
Recalling the estimates of $\partial_{x}^{k+1}F_{1},k=0,1$ in Lemma 4.6 and (4.12), we get
$-\tau\int_{0}^{t}\int_{\mathbb{R}}\partial_{x}^{k}Q\partial_{x}^{k+2}F_{1}\mathrm{d}x\mathrm{d}t$ $=\tau\int_{0}^{t}\int_{\mathbb{R}}\partial_{x}^{k+1}Q\partial_{x}^{k+1}F_{1}\mathrm{d}x\mathrm{d}t$
$\leq C\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}Q\right)^{2}\mathrm{d}x\mathrm{d}t+C\tau^{2}\int_{0}^{t}\|\partial_{x}^{k+1}F_{1}\|_{L^{2}}^{2}\mathrm{d}t$
$\leq C\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}Q\right)^{2}\mathrm{d}x\mathrm{d}t+C\max\{\delta_{1},\delta_{2}\}.$
Therefore, we conclude that
$N_{1}^{k}$ $\leq\nu_{1}\tau\int_{\mathbb{R}}\left(\partial_{x}^{k}Q(t)\right)\mathrm{d}x+C(\nu_{1})\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Psi(t)\right)\mathrm{d}x+C\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}Q\right)^{2}\mathrm{d}x\mathrm{d}t+C\tau\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}\int_{0}^{t}|\dot{X}_{i}|^{2}\mathrm{d}t$
$+C\tau\int_{\mathbb{R}}\left(\partial_{x}^{k}Q_{0}\right)\mathrm{d}x+C\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Psi_{0}\right)\mathrm{d}x+C(\delta_{1}+\delta_{2})\int_{0}^{t}G^{s}\mathrm{d}t+C\max\{\delta_{1},\delta_{2}\}.$
Secondly, for $N_{2}^{k}$, we have
$N_{2}^{k}\leq C\tau^{2}\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}\int_{0}^{t}|\dot{X}_{i}|^{2}\mathrm{d}t+C(\delta_{1}+\delta_{2})\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Psi\right)^{2}\mathrm{d}x\mathrm{d}t.$
For $N_{3}^{k}$, using Young’s inequality, we have
$N_{3}^{k}\leq\frac{\mu}{8}\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k+1}\Psi\right)^{2}\mathrm{d}x\mathrm{d}t+C\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}^{k}Q\right)^{2}\mathrm{d}x\mathrm{d}t.$
$N_{4}^{k}$ can be estimated in the same way as in Lemma 4.6. Specially, for $k=0$, we have
$F_{4}^{0}=-(v-\widetilde{v})\widetilde{\Pi}-\left(\widetilde{v}^{X_{2}}_{2}-v_{m}\right)\widetilde{\Pi}_{1}^{X_{1}}-\left(\widetilde{v}^{X_{1}}_{1}-v_{m}\right)\widetilde{\Pi}_{2}^{X_{2}}.$
So, we have
$N_{4}^{0}\leq C(\delta_{1}+\delta_{2})\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}\Psi\right)^{2}\mathrm{d}x\mathrm{d}t+C(\delta_{1}+\delta_{2})\int_{0}^{t}G^{s}\mathrm{d}t+C\max\{\delta_{1},\delta_{2}\}.$
For $k=1$, the term $F_{3}^{1}$ is the same as in Lemma 4.6. Thus, using similar estimates as in Lemma 4.6, we have
$N_{4}^{1}\leq$ $C(\varepsilon_{1}+\delta_{1}+\delta_{2})\left(\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{xx}\Psi\right)^{2}\mathrm{d}x\mathrm{d}t+\int_{0}^{t}G(U)\mathrm{d}t\right)$
$+C(\delta_{1}+\delta_{2})\int_{0}^{t}\int_{\mathbb{R}}\left(\partial_{x}\Phi\right)^{2}\mathrm{d}x\mathrm{d}t+C\max\{\delta_{1},\delta_{2}\}.$
Therefore, combining the above estimates, we get the desired results. ∎
Next, by using Lemma 4.2, Lemma 4.6, Lemma 4.7 and Lemma 4.8, we are able to prove Proposition 3.1.
Proof of Proposition 3.1: Firstly, multiplying (4.16) and (4.17) by $1+2C_{2}$, $\frac{3}{2}C_{1}+2C_{1}C_{2}$, respectively, we have
$(\frac{1}{4}+C_{2})\int_{0}^{t}\|\partial_{x}\Phi\|_{H^{1}}^{2}\mathrm{d}t+\frac{1}{2}C_{1}\int_{0}^{t}\|\partial_{x}\Psi\|_{H^{1}}^{2}\mathrm{d}t\leq\nu(1+2C_{2})\|\Psi\|_{H^{1}}^{2}+\nu_{1}(\frac{3}{2}C_{1}+2C_{1}C_{2})\tau\|Q\|_{H^{1}}^{2}$ (4.18)
$\qquad+C(\nu_{1})\|\partial_{x}\Psi\|_{H^{1}}^{2}+C(\nu)\|\partial_{x}\Phi\|_{H^{1}}^{2}+C\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}|\dot{X}_{i}(t)|^{2}\mathrm{d}t+C_{3}\int_{0}^{t}\|\partial_{x}Q\|_{H^{1}}^{2}\mathrm{d}t$
$\qquad+C\left(\|\Psi_{0}\|_{H^{2}}^{2}+\|\partial_{x}\Phi_{0}\|_{H^{1}}^{2}+\tau\|Q_{0}\|_{H^{1}}^{2}\right)+C(\varepsilon_{1}+\delta_{1}+\delta_{2})\int_{0}^{t}(G^{s}(U)+G(U))\mathrm{d}t+C\delta_{0},$
where $C_{3}$ is a constant independent of $\tau$ and $T$. Then, by using the smallness of $\delta_{1},\delta_{2}$ and choosing $\nu,\nu_{1}$ to be sufficiently small in equation (4.18), we get
$(\frac{1}{4}+C_{2})\int_{0}^{t}\|\partial_{x}\Phi\|_{H^{1}}^{2}\mathrm{d}t+\frac{1}{2}C_{1}\int_{0}^{t}\|\partial_{x}\Psi\|_{H^{1}}^{2}\mathrm{d}t\leq\frac{1}{2}\|\Psi\|_{L^{2}}^{2}+\frac{1}{2}\tau\|Q\|_{L^{2}}^{2}$ (4.19)
$\qquad+C\left(\|(\partial_{x}\Psi,\partial_{x}\Phi,\sqrt{\tau}\partial_{x}Q)\|_{H^{1}}^{2}+\int_{0}^{t}\|\partial_{x}Q\|_{H^{1}}^{2}\mathrm{d}t\right)+\frac{1}{16}\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}|\dot{X}_{i}(t)|^{2}\mathrm{d}t$
$\qquad+C\left(\|\Psi_{0}\|_{H^{2}}^{2}+\|\partial_{x}\Phi_{0}\|_{H^{1}}^{2}+\tau\|Q_{0}\|_{H^{1}}^{2}\right)+\frac{1}{32}\int_{0}^{t}(G^{s}(U)+G(U))\mathrm{d}t+C\delta_{0}.$
Next, using Lemma 4.6 and choosing $\varepsilon_{1},\delta_{1},\delta_{2}$ sufficiently small, we get that
$(\frac{1}{8}+C_{2})\int_{0}^{t}\|\partial_{x}\Phi\|_{H^{1}}^{2}\mathrm{d}t+\frac{1}{4}C_{1}\int_{0}^{t}\|\partial_{x}\Psi\|_{H^{1}}^{2}\mathrm{d}t\leq\frac{1}{2}\|\Psi\|_{L^{2}}^{2}+\frac{1}{2}\tau\|Q\|_{L^{2}}^{2}+\frac{1}{2}\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}|\dot{X}_{i}(t)|^{2}\mathrm{d}t$ (4.20)
$\qquad+\frac{1}{2}\int_{0}^{t}G(U)\mathrm{d}t+\frac{1}{2}\int_{0}^{t}G^{s}(U)\mathrm{d}t+C\left(\|\Psi_{0}\|_{H^{2}}^{2}+\|\partial_{x}\Phi_{0}\|_{H^{1}}^{2}+\tau\|Q_{0}\|_{H^{2}}^{2}\right)+C\delta_{0}.$
Then, combining the $L^{2}$ estimate (4.3) and (4.20), we get
$\|\Phi\|^{2}_{L^{2}}+\frac{1}{2}\|\Psi\|^{2}_{L^{2}}+\frac{1}{2}\tau\|Q\|^{2}_{L^{2}}+\frac{1}{2}\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}|\dot{X}_{i}|^{2}\mathrm{d}t+\frac{1}{2}\int_{0}^{t}\left(G^{s}(U)+G(U)\right)\mathrm{d}t$ (4.21)
$\qquad\qquad+\frac{1}{8}\int_{0}^{t}\|\partial_{x}\Phi\|_{H^{1}}^{2}\mathrm{d}t+\frac{1}{4}C_{1}\int_{0}^{t}\|\partial_{x}\Psi\|_{H^{1}}^{2}\mathrm{d}t$
$\leq C\left(\|\Psi_{0}\|_{H^{2}}^{2}+\|\Phi_{0}\|_{H^{2}}^{2}+\tau\|Q_{0}\|_{H^{2}}^{2}\right)+C\delta_{0}.$
Thus, combing the high-energy estimates (4.13) and choosing $\varepsilon_{1},\delta_{1},\delta_{2}$ sufficiently small, we derive that
$\|\Phi\|^{2}_{H^{2}}+\|\Psi\|^{2}_{H^{2}}+\tau\|Q\|^{2}_{H^{2}}+\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}|\dot{X}_{i}|^{2}\mathrm{d}t+\int_{0}^{t}\left(G^{s}(U)+\|\left(\Phi_{x},\Psi_{x}\right)\|^{2}_{H^{1}}+\|Q\|^{2}_{H^{2}}\right)\mathrm{d}t$ (4.22)
$\leq C\left(\|\Phi_{0}(\cdot)\|^{2}_{H^{2}}+\|\Psi_{0}(\cdot)\|^{2}_{H^{2}}+\tau\|Q_{0}(\cdot)\|^{2}_{H^{2}}\right)+C\delta_{0}.$
This finish the proof of Proposition 3.1.
5. Proof of Theorem 1.2
In this section, we show the Theorem 1.2 by use of the uniform estimates obtained in Section 4 and usual compactness arguments. Firstly, according to Theorem 1.1, we get
$\sup_{0\leq t<+\infty}\|(\Phi^{\tau},\Psi^{\tau},\sqrt{\tau}Q^{\tau})(t,\cdot)\|_{H^{2}}^{2}+\int_{0}^{+\infty}\left(\|(\Phi^{\tau}_{x},\Psi^{\tau}_{x})\|_{H^{1}}^{2}+\|Q^{\tau}\|_{H^{2}}^{2}\right)\mathrm{d}t\leq C_{0}E(0)+C_{0}\delta_{0},$
where $E(0)=\|(\Phi^{\tau},\Psi^{\tau},\sqrt{\tau}Q^{\tau})(0,\cdot)\|_{H^{2}}$, $\Phi^{\tau}=v^{\tau}-\widetilde{v}^{\tau},\Psi^{\tau}=u^{\tau}-\widetilde{u}^{\tau},Q^{\tau}=\Pi^{\tau}-\widetilde{\Pi}^{\tau}$, $C_{0}$ is a constant independent of $\tau$ and $\widetilde{v}^{\tau}=(\widetilde{v}_{1})^{\tau}+(\widetilde{v}_{2})^{\tau}-v_{m},\widetilde{u}^{\tau}=(\widetilde{u}_{1})^{\tau}+(\widetilde{u}_{1})^{\tau}-u_{m},\widetilde{\Pi}^{\tau}=(\widetilde{\Pi}_{1})^{\tau}+(\widetilde{\Pi}_{2})^{\tau}$ are the compose waves of system (1.3), respectively. Thus, there exist $(\Phi^{0},\Psi^{0})\in L^{\infty}((0,\infty);H^{2})$ and $Q^{0}\in L^{2}((0,\infty);H^{2})$ such that
$(\Phi^{\tau},\Psi^{\tau})\rightharpoonup(\Phi^{0},\Psi^{0})\qquad weakly-*\quad in\quad L^{\infty}((0,\infty);H^{2}),$ (5.1)
$Q^{\tau}\rightharpoonup Q^{0}\qquad weakly-\quad in\quad L^{2}((0,\infty);H^{2}).$
Secondly, from Lemma 2.1, we get
$\|(\widetilde{v}_{1})^{\tau}-v_{m}\|_{L^{2}(\mathbb{R}^{+})}+\|(\widetilde{v}_{1})^{\tau}-v_{-}\|_{L^{2}(\mathbb{R}^{-})}+\|(\widetilde{v}_{1})^{\tau}_{x}\|_{H^{2}}\leq C,$
$\|(\widetilde{v}_{2})^{\tau}-v_{+}\|_{L^{2}(\mathbb{R}^{+})}+\|(\widetilde{v}_{2})^{\tau}-v_{m}\|_{L^{2}(\mathbb{R}^{-})}+\|(\widetilde{v}_{2})^{\tau}_{x}\|_{H^{2}}\leq C,$
$\|(\widetilde{u}_{1})^{\tau}-u_{m}\|_{L^{2}(\mathbb{R}^{+})}+\|(\widetilde{u}_{1})^{\tau}-u_{-}\|_{L^{2}(\mathbb{R}^{-})}+\|(\widetilde{u}_{1})^{\tau}_{x}\|_{H^{2}}\leq C,$
$\|(\widetilde{u}_{2})^{\tau}-u_{+}\|_{L^{2}(\mathbb{R}^{+})}+\|(\widetilde{u}_{2})^{\tau}-u_{m}\|_{L^{2}(\mathbb{R}^{-})}+\|(\widetilde{u}_{2})^{\tau}_{x}\|_{H^{2}}\leq C,$
$\|(\widetilde{\Pi}_{1})^{\tau}\|_{H^{2}}\leq C,\qquad\|(\widetilde{\Pi}_{2})^{\tau}\|_{H^{2}}\leq C,$
where $C$ independent of $\tau$.
Then, using compactness theorem, for $i=1,2$ and any $T>0$, we have
$(\widetilde{v}_{i})^{\tau}\rightarrow(\widetilde{v}_{i})^{0},\quad(\widetilde{u}_{i})^{\tau}\rightarrow(\widetilde{u}_{i})^{0},\qquad strongly\quad in\quad C([0,T];H_{loc}^{2}),$
$(\widetilde{\Pi}_{i})^{\tau}\rightharpoonup(\widetilde{\Pi}_{i})^{0},\qquad weakly-\quad in\quad L^{\infty}((0,\infty),H^{2}).$
In addition, let $\tau\rightarrow 0$ in (2.2), we have $\tau\widetilde{\Pi}_{i}\rightarrow 0$ in $D^{\prime}((0,\infty)\times\mathbb{R})$ and
$(\widetilde{\Pi}_{i})^{\tau}\rightharpoonup\mu\frac{(\widetilde{u}_{i})_{x}^{0}}{(\widetilde{v}_{i})^{0}}:=(\widetilde{\Pi}_{i})^{0},\qquad in\quad\mathcal{D}^{\prime}((0,\infty)\times\mathbb{R})$
and we know that $(\widetilde{v}_{i})^{0},(\widetilde{u}_{i})^{0}$ are the $i$-traveling wave solutions of classical Navier-Stokes equations. Therefore, for any $T>0$, we have
$\widetilde{v}^{\tau}=(\widetilde{v}_{1})^{\tau}+(\widetilde{v}_{2})^{\tau}-v_{m}\rightarrow(\widetilde{v}_{1})^{0}+(\widetilde{v}_{2})^{0}-v_{m}=:\widetilde{v}^{0},\quad strongly\quad in\quad C([0,T];H_{loc}^{2}),$ (5.2)
$\widetilde{u}^{\tau}=(\widetilde{u}_{1})^{\tau}+(\widetilde{u}_{2})^{\tau}-u_{m}\rightarrow(\widetilde{u}_{1})^{0}+(\widetilde{u}_{2})^{0}-u_{m}=:\widetilde{u}^{0},\quad strongly\quad in\quad C([0,T];H_{loc}^{2}),$
$\widetilde{\Pi}^{\tau}=(\widetilde{\Pi}_{1})^{\tau}+(\widetilde{\Pi}_{2})^{\tau}\rightharpoonup(\widetilde{\Pi}_{1})^{0}+(\widetilde{\Pi}_{2})^{0}=:\widetilde{\Pi}^{0},\qquad weakly-\quad in\quad L^{\infty}((0,\infty);H_{loc}^{2}).$
Finally, for any $T>0$, using (4.1), we know that $\Phi_{t}^{\tau}$ and $\Psi_{t}^{\tau}$ are bounded in $L^{2}((0,T);H^{1})$. Furthermore, using compactness theorem, for any $\alpha>0$, $(\Phi^{\tau},\Psi^{\tau})$ are relatively compact in $C([0,T];H_{loc}^{2-\alpha})$. Then, as $\tau\rightarrow 0$, we have
$(\Phi^{\tau},\Psi^{\tau})\rightarrow(\Phi^{0},\Psi^{0})\qquad strongly\quad in\quad C([0,T];H_{loc}^{2-\alpha}).$
Therefore, combining (5.2), we have
$(v^{\tau},u^{\tau})\rightarrow(\Phi^{0}+\widetilde{v}^{0},\Psi^{0}+\widetilde{u}^{0})=:(v^{0},u^{0}),\qquad strongly\quad in\quad C([0,T];H_{loc}^{2-\alpha}).$ (5.3)
On the other hand, noting that $\sqrt{\tau}\Pi^{\tau}$ is uniform bounded in $L^{\infty}((0,\infty);H^{2})$, which yields $\tau\Pi_{t}^{\tau}\rightarrow 0$ in $D^{\prime}((0,\infty)\times\mathbb{R})$ as $\tau\rightarrow 0$. Let $\tau\rightarrow 0$ in (1.3), we have
$\Pi^{\tau}\rightharpoonup\mu\frac{(u^{0})_{x}}{v^{0}}:=\Pi^{0}\qquad a.e.\quad(0,\infty)\times\mathbb{R}$ (5.4)
and we conclude that $v^{0},u^{0}$ are the solutions of classical Navier-Stokes equations. Then, combining (5.1), (5.2), (5.3) and (5.4), we get the desired results.
Appendix A proof of Lemma 2.2
Firstly, we define $F_{i}(t,X_{i})$ ($i=1,2$) as follows:
$F_{i}(t,X_{i})=-\frac{M}{\delta_{i}}\left[\int_{\mathbb{R}}\frac{a}{\sigma_{i}}(\widetilde{u}_{i}^{X_{i}})_{x}(p(v)-p(\widetilde{v}))\mathrm{d}x-\int_{\mathbb{R}}a\left(p(\widetilde{v}_{i}^{X_{i}})\right)_{x}(v-\widetilde{v})\mathrm{d}x\right].$
Applying (2.14), Lemma 2.1 and the definition of $a(t,x)$ in (2.11), we can obtain that
$\|a\|_{C^{1}}\leq 2,\quad\|\widetilde{v}_{i}^{X_{i}}\|_{C^{2}}<\infty,\quad\|(\widetilde{v}_{i}^{X_{i}})_{x}\|_{W^{1,1}}\leq C\delta_{i},\quad\|(\widetilde{u}_{i}^{X_{i}})_{x}\|_{W^{1,1}}\leq C\delta_{i}.$
Then, we have
$\sup\limits_{X_{i}\in\mathbb{R}}|F_{i}(t,X_{i})|$ $\leq\frac{C}{\delta_{i}}\|a\|_{C^{1}}\||p(\widetilde{v}_{i})|,|p(v)|,|\widetilde{v}_{i}|,|v|\|_{L^{\infty}}\left(\|(\widetilde{v}_{i}^{X_{i}})_{x}\|_{L^{1}}+\|(\widetilde{u}_{i}^{X_{i}})_{x}\|_{L^{1}}\right)$ (A.1)
$\leq C,$
and
$\sup\limits_{X_{i}\in\mathbb{R}}|\partial_{X_{i}}F_{i}(t,X_{i})|$ $\leq\frac{C}{\delta_{i}}\|a\|_{C^{1}}\||p(\widetilde{v}_{i})|,|p(v)|,|\widetilde{v}_{i}|,|v|\|_{L^{\infty}}\left(\|(\widetilde{v}_{i}^{X_{i}})_{xx}\|_{L^{1}}+\|(\widetilde{u}_{i}^{X_{i}})_{xx}\|_{L^{1}}\right)$
$\leq C,$
where $C$ is a constant independent of $t$ and $\tau$. Therefore, the ODE (2.10) has a unique Lipschitz continuous solution by using the Cauchy-Lipschitz theorem (Lemma C.1 of [22]).
In particularly, since $|\dot{X}_{1}(t)|+|\dot{X}_{2}(t)|\leq C$ according (A.1), we have (2.15).
Appendix B proof of Lemma 4.3
Although the proof of Lemma 4.3 can be directly derived from the Appendix A of [10], there are still slight differences between the traveling wave solutions of the relaxed system and classical system. Therefore, we still provide the proof of the Lemma 4.3 in this appendix. We only estimates the case of $i=2$, other case $i=1$ can be followed in a similar way. Firstly, according Lemma 2.1, we can obtain that
$|(\widetilde{v}_{2})^{X_{2}}_{x}|\leq C\delta_{2}^{2}e^{-C\delta_{2}|x-\sigma_{2}t-X_{2}(t)|}\quad\forall x\in\mathbb{R},\quad t>0,$
and
$|\widetilde{v}-(\widetilde{v}_{2})^{X_{2}}|=|(\widetilde{v}_{1})^{X_{1}}-v_{m}|\leq\begin{cases}C\delta_{1}e^{-C\delta_{1}|x-\sigma_{1}t-X_{1}(t)|}\quad\text{if}\quad x\geq\sigma_{1}t+X_{1}(t),\\ C\delta_{1}\quad\quad\quad\qquad\qquad\qquad\text{if}\quad x\leq\sigma_{1}t+X_{1}(t).\end{cases}$
Noting that $X_{1}(t)+\sigma_{1}t\leq\frac{\sigma_{1}t}{2}\leq 0\leq\frac{\sigma_{2}t}{2}\leq X_{2}(t)+\sigma_{2}t$, we can derive that
$|(\widetilde{v}_{2})^{X_{2}}_{x}||(\widetilde{v}_{1})^{X_{1}}-v_{m}|\leq\begin{cases}C\delta_{1}\delta_{2}^{2}e^{-C\delta_{1}|x-\sigma_{1}t-X_{1}(t)|}\quad\text{if}\quad x\geq 0,\\ C\delta_{1}\delta_{2}^{2}e^{-C\delta_{2}|x-\sigma_{2}t-X_{2}(t)|}\quad\text{if}\quad x\leq 0.\end{cases}$
On the other hand, we note that
$x-\sigma_{1}t-X_{1}(t)\geq-\frac{\sigma_{1}t}{2}\geq 0\quad\text{if}\quad x\geq 0,$
$x-\sigma_{2}t-X_{2}(t)\leq-\frac{\sigma_{2}t}{2}\leq 0\quad\text{if}\quad x\geq 0.$
Therefore, we conclude that
$|(\widetilde{v}_{2})^{X_{2}}_{x}||(\widetilde{v}_{1})^{X_{1}}-v_{m}|\leq C\delta_{1}\delta_{2}^{2}e^{-C\min\{\delta_{1},\delta_{2}\}t},\quad\forall x\in\mathbb{R},\quad t>0,$
and
$\int_{\mathbb{R}}|(\widetilde{v}_{2})^{X_{2}}_{x}||(\widetilde{v}_{1})^{X_{1}}-v_{m}|\mathrm{d}x\leq C\delta_{1}\delta_{2}^{2}e^{-C\min\{\delta_{1},\delta_{2}\}t},\quad t>0.$
In a similar way, using Lemma 2.1, we can get
$|(\widetilde{v}_{1})^{X_{1}}_{x}||(\widetilde{v}_{2})^{X_{2}}_{x}|\leq C\delta_{1}^{2}\delta_{2}^{2}e^{-C\min\{\delta_{1},\delta_{2}\}t},\quad\forall x\in\mathbb{R},\quad t>0,$
and
$\int_{\mathbb{R}}|(\widetilde{v}_{1})^{X_{1}}_{x}||(\widetilde{v}_{2})^{X_{2}}_{x}|\mathrm{d}x\leq C\delta_{1}\delta_{2}e^{-C\min\{\delta_{1},\delta_{2}\}t},\quad t>0.$
Acknowledgement: Yuxi Hu’s Research is supported by the Fundamental Research Funds for the Central Universities (No. 2023ZKPYLX01) .
Data Availibility: Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Conflict of interest: On behalf of all authors, the corresponding author states that there is no conflict of interest.